### v1

Abstract
Decision trees are a ubiquitous model for classification and regression tasks due to their interpretability and efficiency. However, solving the optimal decision tree (ODT) problem remains a challenging combinatorial optimization task. Even for the simplest splitting rules—axis-parallel hyperplanes—it is NP-hard to optimize. In Part I of this series, we rigorously defined the proper decision tree model through four axioms and, based on these, introduced four formal definitions of the ODT problem. From these definitions, we derived four generic algorithms capable of solving ODT problems for arbitrary decision trees satisfying the axioms. We also analyzed the combinatorial geometric properties of hypersurfaces, showing that decision trees defined by polynomial hypersurface splitting rules satisfy the proper axioms that we proposed.
In this second paper (Part II) of this two-part series, building on the algorithmic and geometric foundations established in Part I, we introduce the first hypersurface decision tree (HODT) algorithm. To the best of our knowledge, existing optimal decision tree methods are, to date, limited to hyperplane splitting rules—a special case of hypersurfaces—and rely on general-purpose solvers. In contrast, our HODT algorithm addresses the general hypersurface decision tree model without requiring external solvers.
In experiments, we provide a comprehensive empirical analysis of the combinatorial complexity of hyperplane decision trees. We implemented the two most suitable algorithms from Part I and compared their performance in both sequential and parallelized settings. We then selected the best-performing method and evaluated its generalization on both synthetic and real-world datasets. Using synthetic datasets generated from ground-truth hyperplane decision trees, we vary tree size, data size, dimensionality, and label and feature noise. Results showing that our algorithm recovers the ground truth more accurately than axis-parallel trees and exhibits greater robustness to noise. We also analyzed generalization performance across 30 real-world datasets, showing that HODT can achieve up to 30% higher accuracy than the state-of-the-art optimal axis-parallel decision tree algorithm when tree complexity is properly controlled.
(a) Ground truth, 0-1 loss: 0, KK=5, dd=3
(b) CART, 0-1 loss: 56, KK=15, dd=5
(c) ConTree, 0-1 loss: 34, KK=15, dd=4
(d) HODT (Our), 0-1 loss: 18, KK=3, dd=3
Figure 1: Synthetic dataset and decision trees with corresponding 0-1 loss, tree sizes ($K$), and tree depths ($d$).
1 Introduction
A decision tree is a supervised machine learning model to which makes predictions using a tree-based subdivision of the feature space. Imagine a flowchart or a series of “yes” or “no” questions that guide towards a final decision. Owing to their simplicity and interpretability, decision trees and their variants—such as classification and regression trees (CART) (Breiman et al., 1984), C4.5 (Quinlan, 2014) and random forests (Breiman, 2001)—have achieved wide success across diverse fields. Beyond interpretability, decision tree methods are also highly accurate: empirical studies have shown that tree-based algorithms often outperform more complex deep learning classifiers on tabular datasets, where data are structured with meaningful features (Grinsztajn et al., 2022; Shwartz-Ziv and Armon, 2022).
Nevertheless, the classical axis-parallel decision tree (ADT) has important limitations. Geometrically, each internal node partitions the feature space by testing a single feature value, producing a split parallel to one coordinate axis. For example, a node may ask:“ Is feature $x_{d}$ greater than some value vv?” Such a splitting rule divides the space into two half-spaces, $x_{d}\leq v$ and $x_{d}>v$, separated by a hyperplane parallel to the $x_{d}=0$. Consequently, ADTs can represent only hyper-rectangular decision regions. This restriction forces axis-parallel trees to approximate complex boundaries with a “staircase-like” structure, as illustrated by Murthy et al. (1994).
By contrast, decision trees that allow hyperplane splits are often both smaller and more accurate. When the underlying decision regions involve complex or non-convex boundaries, trees with hyperplane or hypersurface splitting rules provide a more faithful representation, whereas axis-parallel trees remain limited to hyper-rectangular partitions. For instance, Figure LABEL:fig:_comparision shows a synthetic dataset generated from three quadratic hypersurfaces (conic sections). The decision regions produced by classical CART and by the state-of-the-art optimal decision tree algorithm ConTree (Brița et al., 2025), displayed in the two middle panels, are constrained to irregular axis-aligned rectangles. Both CART and ConTree yield trees with 15 splitting rules, of depths 5 and 4, respectively.
The above example motivates the design of algorithms for constructing decision trees with more flexible splitting rules, such as hyperplanes or, more generally, hypersurfaces. However, to the best of our knowledge, no algorithm—whether exact, approximate, or heuristic—has been proposed for learning decision trees with hypersurface splitting rules. In this second paper of the series, we present the first such algorithm, building on the results from Part I to solve the size-constrained optimal hypersurface decision tree (HODT) problem. The algorithm is developed through several key results, including the prefix-closed properties of crossed hyperplanes and the incremental generation of ancestry relation matrices. The main contributions of this paper are as follows:
  • Crossed-hyperplane property: We identify an important ancestry relations for pair of hyperplanes, importantly we proved that proper decision tree can not be constructed from a combination of decision trees that contains a pair of crossed hyperplanes, which substantially reduce the search space of the HODT problem.
  • Incremental ancestry-matrix algorithm: Exploiting the prefix-closed property of crossed hyperplanes, and incremental evaluation property of ancestry relation matrix, we develop an efficient incremental algorithm for generating the ancestry relation matrix.
  • Application of size-constrained ODT algorithms to HODT problem: We extend the two rigorously proven algorithms from Part I to solve the size-constrained HODT problem, covering classical axis-parallel hyperplanes, general hyperplanes, and hypersurfaces. Our algorithms run in polynomial time (for fixed tree size and feature dimension), are embarrassingly parallelizable (parallelizable without communication).
  • Synthetic data experiments: Using two proposed heuristic methods, we conduct extensive experiments on hyperplane decision algorithms with synthetic datasets generated by hyperplane decision trees (polygonal regions). Results show that hyperplane decision trees recover the ground truth more accurately than both heuristics (size-/depth-constrained CART) and Brița et al. (2025)’s optimal axis-parallel tree algorithm (ConTree) under various noise conditions. Remarkably, our algorithms sometimes surpass the ground truth accuracy on noisy data but still out performs well in out-of-sample test, countering the notion that optimal algorithms necessarily overfit.
  • Benchmarking on real-world datasets: We evaluate performance on 30 real-world datasets, comparing out algorithms with CART and ConTree. The results demonstrate that hypersurface-based decision trees achieve superior generalization accuracy when tree size is appropriately controlled.
The structure of this series is as follows. Part I introduced the algorithmic and geometric foundations of the ODT problem. Part II, presented here, develops the first algorithm for solving the HODT problem, building directly on the foundations of Part I. Section 3, details the construction of the algorithm: Subsection 3.1 introduces ancestry relations for hyperplanes, identifying a critical property—crossed hyperplanes—and proving that any set of hyperplanes containing such a pair cannot form a proper decision tree. This result forms the basis for an efficient ancestry-matrix generator free of crossed hyperplanes, described in Subsection 3.2. The complete HODT algorithm is then presented in Subsection 3.3.
Section 4 addresses the combinatorial intractability of solving the HODT problem by introducing two heuristic methods: $\mathit{hodtCoreset}$ ($\mathit{hodt}$ over coreset) and $\mathit{hodtWSH}$ (HODT with selected hyperplanes), which yield practical solutions for large-scale experiments.
Finally, Section 5, presents a detailed runtime analysis of the algorithms $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ under both sequential and parallel execution. The better-performing method is then used for empirical evaluation of generalization performance on synthetic datasets (5.3) and real-world datasets (Subsection 5.4).
2 Review of Part I
General framework for solving the hypersurface decision tree problem
In Part I, we formalized the decision tree problem through a set of axioms, referred to as the proper decision tree axiom. Based on these axioms, we derived four generic definitions of the decision tree problem, expressed unambiguously as four recursive programs. Through straightforward derivations, we proved both the existence and non-existence of dynamic programming (DP) solutions for these definitions under certain objective functions. We then examined the feasibility of solving the ODT problem under depth or size constraints, and argued that size-constrained trees are more suitable for problems with formidable combinatorial complexity. This suitability arises from their inherent parallelizability, achieved by factoring size-constrained trees via $K$-combinations.
To recap, given a list of data $\mathit{xs}:\left[\mathbb{R}^{D}\right]$ and splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$ that satisfy the proper decision tree axiom, the size-constrained ODT problem can be solved by the following program:
$\mathit{odt}_{\text{size}}\left(K\right)=\mathit{min}_{E}\circ\mathit{concatMapL}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}_{K}$ (1)
where $\mathit{kcombs}\left(K\right):\left[\mathcal{R}\right]\to\left[\left[\mathcal{R}\right]\right]$ takes a list of rules and returns all possible $K$-combinations. The program $\mathit{sodt}$, a DP algorithm, solves a simplified decision tree problem for each $K$-combination using $\mathit{concatMapL}_{\mathit{sodt}\left(\mathit{xs}\right)}$. The optimal decision tree is then selected by $\mathit{min}_{E}$ given some objective $E$.
In particular, we provide three definitions of $\mathit{sodt}$: $\mathit{sodt}_{\text{rec}}$, $\mathit{sodt}_{\text{vec}}$, and $\mathit{sodt}_{\text{kperms}}$. It remains unclear which of these three formulations is most effective for size-constrained ODT algorithms. The recursive definition $\mathit{sodt}_{\text{rec}}$ requires provably fewer operations than the others and is defined as a DP algorithm:
$\mathit{sodt}_{\text{rec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)$ (2)
$\mathit{sodt}_{\text{rec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{sodt}_{\text{rec}}$ $\left(\mathit{xs},\left[r\right]\right)=\left[\mathit{DN}\left(\mathit{DL}\left(xs^{+}\right),r,\mathit{DL}\left(xs^{-}\right)\right)\right]$
[…]
Alternatively, $\mathit{sodt}_{\text{vec}}$ is defined as
$\mathit{sodt}_{\text{vec}}=\mathit{min}_{E}\circ\mathit{genDTs}_{\text{vec}}$ (3)
where $\mathit{genDTs}_{\text{vec}}$ is defined sequentially, with each recursive step processing only one rule $r$:
$\mathit{genDTs}_{\text{vec}}$ $:\mathcal{D}\times\left[\mathcal{R}\right]\to\left[\mathit{DTree}\left(\mathcal{R},\mathcal{D}\right)\right]$ (4)
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right]$
The definition in 3 is a two-phase process: $\mathit{min}_{E}$ is applied to $\mathit{genDTs}_{\text{vec}}$ only after all configurations are exhaustively generated. This leads to significantly higher computational cost compared with $\mathit{sodt}_{\text{rec}}$, which integrates $\mathit{min}_{E}$ directly into the recursion. We further prove that enforcing $\mathit{min}_{E}$ directly inside $\mathit{genDTs}_{\text{vec}}$ results in non-optimal solutions.
Since $\mathit{odt}_{\text{size}}$ can be instantiated with any of the three definitions of $\mathit{sodt}$, a choice must be made in practice. As discussed in Part I, $\mathit{sodt}_{\text{rec}}$ has stronger theoretical efficiency due to its DP construction. However, the performance of an algorithm often depends not only on its theoretical design but also on the hardware used for implementation. In this regard, the sequential nature of $\mathit{sodt}_{\text{vec}}$, and $\mathit{sodt}_{\text{kperms}}$ makes them particularly amenable to vectorized operations and parallel execution. Modern CPUs and GPUs are highly optimized for parallel vectorized computations, whereas the recursive structure of $\mathit{sodt}_{\text{rec}}$ hinders full vectorization when processing data in batches.
In this paper, we conduct a detailed computational analysis of these two approaches under both parallel and sequential settings. The better-performing method is then employed in the final experiments to evaluate generalization performance.
Optimal hypersurface decision tree problem
In recalling the general framework for solving the size-constrained ODT problem, an astute reader may notice that the program $\mathit{odt}_{\text{size}}$ operates on a list of splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$ rather than directly on the data. This design choice reflects our aim of providing a generic, simple, and modular framework, since decision trees are defined in terms of splitting rules rather than data instances. Consequently, when applying the framework to practical ODT problems in machine learning, a separate procedure $\mathit{gen}_{\mathit{splits}}$ must be introduced to generate the splitting rules, with its definition depending on the specific task. This modularity also provides a further advantage: adapting the framework to different problem settings requires only minimal code modifications.
Formally, the complete algorithm for solving the HODT problem is given by $\mathit{hodt}^{M}:\mathbb{N}\times\mathcal{D}\to\mathit{DTree}\left(\mathcal{H}^{M},\mathcal{D}\right)$, which can be specified as
$\mathit{hodt}_{K}^{M}$ $=\mathit{odk}_{\text{size}}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$ (5)
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M}$
where $\mathit{embed}_{M}:\left[\mathbb{R}^{D}\right]\to\left[\mathbb{R}^{G}\right]$ is the Veronese embedding ($\rho_{M}$) introduced in Part I, mapping data points $\mathit{xs}$ in $\mathbb{R}^{D}$ to the embedded dataset $\rho_{M}\left(\mathit{xs}\right)$ in $\mathbb{R}^{G}$, with $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$. Moreover, $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ can be unified into a single program $\mathit{nestedCombs}\left(K,G\right)$, for which the implementation was given in Part I. For completeness, we include its pseudocode in Algorithm 5 in Appendix A.
However, the specification (5) is not directly executable. Specifically, it makes a hidden assumption regarding $\mathit{sodt}$ that the ancestry relation matrices $\boldsymbol{K}$ are precomputed and available in memory with $O\left(1\right)$ access. In practice, these matrices must be generated for each $K$-combination of splitting rules before applying $\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}$. Thus, an algorithm is required to efficiently generate these matrices.
To address this issue, we want to modify $\mathit{nestedCombs}\left(K,D\right)$ so that it directly incorporates ancestry relation matrix computation. In particular, the new generator should satisfy two properties:
  1. 1.
    It filters out all nested combinations containing crossed hyperplanes.
  2. 2.
    It computes the ancestry relation matrix for each valid nested combination.
We denote this new generator by $\mathit{nestedCombsFA}$, short for “nested combination generator with filtering and ancestry updates.” Formally,
$\mathit{\mathit{nestedCombsFA}}\left(K,G\right)=\mathit{filter}_{p}\circ\mathit{mapL}_{\mathit{calARMs}}\circ\mathit{nestedCombs}\left(K,G\right)$ (6)
where $\mathit{calARM}:\mathit{NC}\to\mathit{NCR}$ computes the ancestry relation matrix of a $\left(K,G\right)$-nested combinations, and the predicate $p$ checks whether a combination is free of crossed hyperplanes. Since the ancestry relation matrix $\boldsymbol{K}$ for a $K$-length rule list $\mathit{rs}_{K}$ is a $K\times K$ square matrix. Both $\boldsymbol{K}$ and $\mathit{rs}_{K}$ are integer-valued and can be stored compactly by stacking $\mathit{rs}_{K}$ above $\boldsymbol{K}$. That is, a configuration $\mathit{ncr}:\mathit{NCR}$ is defined as
$\mathit{ncr}=\left[\begin{array}[]{c}\mathit{rs}_{K}\\ \boldsymbol{K}\end{array}\right]$ (7)
Thus, a collection of $M$ configurations can be stored as a tensor of size $M\times\left(K+1\right)\times K$.
As usual, post-hoc computation of ancestry matrices after (6) is inefficient. A natural question arises, echoing similar discussions in Part I: can we fuse $\mathit{filter}_{p}$ and $\mathit{calARM}$ directly into the recursive definition of $\mathit{\mathit{nestedCombs}}$? The answer is yes. In the next section, we show how to construct such a fused generator, yielding an efficient program that simultaneously generates feasible nested combinations and their associated ancestry relation matrices within a single recursive procedure.
3 Optimal hypersurface decision tree algorithm
3.1 Ancestry relation and crossed hyperplanes
According to the proper decision tree axioms, descendant rules can only be generated from decision regions determined by their ancestor rules (Axioms 1 and 2). For any pair of splitting rules $r_{i}$ and $r_{j}$ with $i\neq j$, exactly one of the following holds:: $r_{i}\swarrow r_{j}$, $r_{i}\searrow r_{j}$, and $r_{i}\overline{\left(\swarrow\vee\searrow\right)}r_{j}$. In this section, we focus on constructing the ancestry relations $h_{i}^{M}\swarrow h_{j}^{M}$, $h_{i}^{M}\searrow h_{j}^{M}$, and $h_{i}^{M}\overline{\left(\swarrow\vee\searrow\right)}h_{j}^{M}$ for hypersurface splitting rules $h_{i}^{M}$ and $h_{j}^{M}$, defined by degree $M$ polynomials, which lays out foundations for discussing how to construct the ancestry relation matrix $\boldsymbol{K}$ efficiently for each $K$-combination of hyperplanes prior to executing $\mathit{sodt}$. Since we have already established the equivalence between polynomial hypersurface classification and hyperplane classification in the embedded space, we restrict our discussion here to the case $M=1$ (i.e., hyperplanes). The general results for higher-degree hypersurfaces follow directly as the result of Theorem 11 in Part I.
Ancestry relation between hyperplanes in general position
Axiom 4 of proper decision trees is highly problem-dependent. For axis-parallel decision trees, the splitting rule $h^{0}$ can be characterized by a single data item. Assuming no two data points lie on the same axis-parallel hyperplane (i.e., the set of candidate splitting rules contains no duplicates—duplicates, if present, can simply be removed without loss of generality), it follows that the relation $h_{i}^{0}\overline{\left(\swarrow\vee\searrow\right)}h_{j}^{0}$ is impossible. This is because any data point always lies on exactly one side of an axis-parallel hyperplane. Thus, for any pair of axis-parallel hyperplanes $h_{i}^{0}$ and $h_{j}^{0}$ either $h_{i}^{0}\swarrow h_{j}^{0}$ or $h_{i}^{0}\searrow h_{j}^{0}$ holds. In other words, any axis-parallel hyperplane $h_{j}^{0}$ must lie in either the left or right subregion determined by another axis-parallel hyperplane $h_{i}^{0}$.
In contrast, for general hyperplanes characterized by data points in general position, ancestry relations are more nuanced. The following fact highlights the key observation about the possible ancestry relations between any pair of such hyperplanes.
Fact 1.
Let $h_{i}$ and $h_{j}$ be two hyperplanes defined by $D$ data points in general position. Their ancestry relation can fall into one of three categories:
  1. 1.
    Mutual ancestry: Both $h_{i}$ and $h_{j}$ can serve as ancestors of each other; that is, both $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ and $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ are viable.
  2. 2.
    Asymmetrical ancestry: Only one hyperplane can be the ancestor of the other; that is, exactly one of $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ or $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ holds.
  3. 3.
    No ancestry: Neither hyperplane can be the ancestor of the other; in this case, both $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ or $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ fail to hold. Here, $h_{i}$ and $h_{j}$ are said to cross each other, and we refer to them as a pair of crossed hyperplanes.
Figure 2: Three possible ancestry relations between two hyperplanes in $\mathbb{R}^{2}$, the black black circles represent data points used to define these hyperplanes.
Figure 3: Three equivalent representations describing the ancestral relations between hyperplanes. A 44-combination of lines (left), each defined by two data points (black points) in $\mathbb{R}^{2}$, where black arrows represent the normal vectors to the corresponding hyperplanes. The ancestry relation graph (middle) captures all ancestry relations between hyperplanes. In this graph, nodes represent hyperplanes, and arrows represent ancestral relations. An incoming arrow to a node $h_{i}$ indicates that the defining data of the corresponding hyperplane lies on the negative side of $h_{i}$. The absence of an arrow indicates no ancestral relation. Outgoing arrows represent hyperplanes whose defining data lies on the positive side of $h_{i}$. The ancestral relation matrix (right) $\boldsymbol{K}$, where the elements $\boldsymbol{K}_{ij}=1$, $\boldsymbol{K}_{ij}=-1$, and $\boldsymbol{K}_{ij}=0$ indicate that $h_{j}$ lies on the positive side, negative side of $h_{i}$, or that there is no ancestry relation between them, respectively.
Figure 2 illustrates examples of the three ancestry cases for hyperplanes in $\mathbb{R}^{2}$. Figure 3 depicts the corresponding ancestry relations and the ancestry relation graph/matrix for four hyperplanes in $\mathbb{R}^{D}$. The third case—no ancestry relation—requires further discussion. In such cases, constructing a valid decision tree appears impossible, as neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ or $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable. To formalize this intuition, we define the concept of crossed hyperplane formally.
Definition 1.
Crossed hyperplane. A hyperplane $h$ defined by a normal vector $\boldsymbol{w}\in\mathbb{R}^{D+1}$ naturally partitions the space into two disjoint and continuous regions: $h^{+}=\left\{x:\mathbb{R}^{D}\mid\boldsymbol{w}^{T}\bar{x}\geq 0\right\}$ and $h^{-}=\left\{x:\mathbb{R}^{D}\mid\boldsymbol{w}_{i}^{T}\bar{x}<0\right\}$. A pair of hyperplanes $h_{i}$ and $h_{j}$, defined by two sets of $D$ data points $\mathit{xs}$ and $\mathit{ys}$, is said to be crossed, denoted using predicate $p_{\text{crs}}\left(h_{i},h_{j}\right)=\text{True}$ if there exist points $x,x^{\prime}\in\mathit{xs}$ and $y,y^{\prime}\in\mathit{ys}$ such that
$\left(x\in h_{j}^{+}\wedge x^{\prime}\in h_{j}^{-}\right)\wedge\left(y\in h_{i}^{+}\wedge y^{\prime}\in h_{i}^{-}\right).$ (8)
Intuitively, $p_{\text{crs}}\left(h_{i},h_{j}\right)=\text{True}$ if at least two points defining $h_{i}$ lie on opposite sides of $h_{j}$, and simultaneously, at least two points defining $h_{j}$ lie on opposite sides of $h_{i}$.
We hypothesize that any set of rules containing a crossed hyperplane cannot form a proper decision tree. To verify this rigorously, one might consider the existence of a “third hyperplane” $h_{k}$ that separates $h_{i}$ and $h_{j}$ into different branches. If such a hyperplane exists $h_{i}$ and $h_{j}$ would not require an ancestry relation, as they would reside on different branches of $h_{k}$.
The following theorem confirms that no such hyperplanes $h_{k}$ can exist when $h_{i}$ and $h_{j}$ are crossed. Consequently, any combination of hyperplanes containing a crossed pair cannot yield a proper decision tree.
Theorem 1.
If two hyperplanes $h_{i}$ and $h_{j}$ cross each other then: no ancestry relation exists between $h_{i}$ and $h_{j}$, and no hyperplanes $h_{k}$ can separate $h_{i}$ and $h_{j}$ into different branches. Consequently, any combination of hypersurfaces containing such crossed hypersurfaces cannot form a proper decision tree.
Proof.
We prove this by contradiction. Assume there exists a third hypersurface $h_{k}$ such that $h_{i}$ and $h_{j}$ lie on the left and right branches of $h_{k}$, respectively. Then $h_{k}$ must also separate the data points defining $h_{i}$ and $h_{j}$ in two decision regions. By the definition of convex combinations in Euclidean space, for any set of points $xs=\left\{\boldsymbol{x}_{1},\boldsymbol{x}_{2}\ldots\boldsymbol{x}_{D}\right\}$ on the same side of a hyperplane, i.e., $\boldsymbol{w}^{T}\boldsymbol{x}_{i}\geq 0$ or $\boldsymbol{w}^{T}\boldsymbol{x}_{i}<0$, $\forall x_{i}\in xs$, the convex combination $\boldsymbol{y}=\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}$ will satisfy $\boldsymbol{w}^{T}\boldsymbol{y}=\boldsymbol{w}^{T}\sum_{i=1}^{m}\lambda_{i}\boldsymbol{x}_{i}\geq 0$, where $\lambda_{i}\geq 0$. Thus, the convex hull of points on one side of a hyperplane remains entirely on that side.
Therefore, if $h_{k}$ separate the defining data points of $h_{i}$ and $h_{j}$ it must also separate their respective convex hulls. However, by the definition of crossed hypersurfaces, the defining points of $h_{i}$ lies on both side of $h_{j}$, and the defining points of $h_{j}$ lies on both sides of $h_{i}$. Consequently, the convex hulls of $h_{i}$ and $h_{j}$ intersect, making it impossible for a single hyperplane $h_{k}$ to separate them. This contradicts the assumption that such $h_{k}$ exists. Hence, no hypersurface can separate crossed hypersurfaces, and any combination containing them cannot form a proper decision tree. ∎
Figure 4: An example illustrating the proof of Fact 1. Demonstrating that the data items $a$, $b$, which define $h_{i}$, and $c$, $d$, which define $h_{j}$ cannot be classified into the disjoint regions defined by a third hyperplane (red).
Again, the conclusion for hypersurfaces follows directly by considering the embedding of the dataset into a higher-dimensional space. An example of the hyperplane case in $\mathbb{R}^{2}$ is shown in Figure 4. The convex hull formed by points $a$, $b$ (which defines $h_{i}$) intersects the convex hull formed by points $c$, $d$ (defining $h_{j}$) at point $e$. Any third hyperplane in $\mathbb{R}^{D}$, such as the red line in the figure, cannot separate the points $a,b$ and $c,d$ into distinct regions.
Lemma 1.
Given a list of $K$ hyperplanes $\mathit{hs}_{K}=\left[h_{1},h_{2},\ldots,h_{K}\right]$ with ancestry matrix $\boldsymbol{K}$. If $\mathit{hs}_{K}$ contains a pair of crossed hyperplanes $h_{i}$ and $h_{j}$, then $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$. If $\mathit{hs}_{K}$ contains no crossed hyperplanes, it can form at least one proper decision tree.
Proof.
By definition, if a pair $h_{i}$ and $h_{j}$ in $\mathit{hs}_{K}$ are crossed, then neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ nor $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable, and therefore $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$. Conversely, if $\mathit{hs}_{K}$ contains no crossed hyperplanes, then for any pair $h_{i},h_{j}\in\mathit{hs}_{K}$, either a mutual ancestry or an asymmetrical ancestry exists. Hence, a proper decision tree can be constructed. ∎
A combination of hyperplanes $\mathit{hs}_{K}$ without crossed hyperplanes is referred to as a feasible combination. Since hyperplanes themselves are nested combinations, we may also refer to them as feasible nested combinations.
Lemma 1, although intuitively obvious, is crucial for constructing an efficient algorithm for generating a crossed-hyperplane-free ancestry matrix $\boldsymbol{K}$, It provides two important computational advantages:
  1. 1.
    To detect crossed hyperplanes, it suffices to check whether symmetric elements across the diagonal of the matrix are zero, which is computationally inexpensive.
  2. 2.
    Only feasible combinations need to be considered; infeasible combinations can be filtered out, reducing computational overhead.
Although the theoretical number of $K$-combinations of hyperplanes for $N$ data points in general position is $O\left(N^{DK}\right)$, Theorem 1 implies that, empirically, the number of feasible combinations is much smaller than this upper bound (see Subsection 5.2). Furthermore, as $K$ increases, we observe empirically that the number of feasible nested combinations decreases once a certain threshold is reached.
3.2 A generic, incremental, crossed-hyperplane free, ancestry relation matrix generator
Incremental update of the ancestry relation matrix
A critical observation is that the ancestry relation matrix can be updated incrementally. Specifically, consider a partial configuration $\mathit{ncr}_{K-1}$ consisting of a list of rules $\mathit{rs}_{K-1}$ and a $\left(K-1\right)\times\left(K-1\right)$ ancestry relation matrix $\boldsymbol{K}^{\prime}$. We can extend $\mathit{ncr}_{K-1}$ to a complete configuration $\mathit{ncr}_{K}$ as follows: first, appending a new rule $r$ to $\mathit{rs}_{K-1}$; second, Compute the ancestry relation of $r$ with each rule in $\mathit{rs}_{K-1}$, and vice versa. This yields the following incremental (sequentially recursive) program $\mathit{calARM}$ for constructing $\mathit{ncr}_{K}$
[…]
$\mathit{calARM}$ $\left(r:\mathit{rs}\right)=\mathit{update}_{\text{arMat}}\left(r,\mathit{calARM}\left(\mathit{rs}\right)\right)$
where $\left[\begin{array}[]{c}r\\ 0\end{array}\right]$ represents a $2\times 1$ matrix. Let $\mathit{calARM}\left(\mathit{rs}\right)=\left[\begin{array}[]{c}\mathit{rs}_{k}\\ \boldsymbol{K}^{\prime}\end{array}\right]$, the update function $\mathit{update}_{\text{arMat}}:\mathit{NCR}\to\mathit{NCR}$ is defined as
$\begin{aligned} \mathit{update}_{\text{arMat}}&\left(r_{j},\left[\begin{array}[]{c}\mathit{rs}_{k}\\ \boldsymbol{K}^{\prime}\end{array}\right]\right)=&\left[\begin{array}[]{ccccc}r_{0}&r_{1}&\ldots&r_{k-1}&r_{j}\\ \boldsymbol{K}_{0,0}^{\prime}&\boldsymbol{K}_{01}^{\prime}&\ldots&\boldsymbol{K}_{0,k-1}^{\prime}&\mathit{AR}_{0,j}\\ \boldsymbol{K}_{1,0}^{\prime}&\boldsymbol{K}_{1,1}^{\prime}&\ldots&\boldsymbol{K}_{1,k-1}^{\prime}&\mathit{AR}_{1,j}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \boldsymbol{K}_{k-1,0}^{\prime}&\boldsymbol{K}_{k-1,1}^{\prime}&\ldots&\boldsymbol{K}_{k-1,k-1}^{\prime}&\mathit{AR}_{k-1,j}\\ \mathit{AR}_{j,0}&\mathit{AR}_{j,1}&\ldots&\mathit{AR}_{j,k-1}&0\end{array}\right]\end{aligned},$ (9)
which extends a $\left(k+1\right)\times k$ matrix to a $\left(k+2\right)\times\left(k+1\right)$ matrix. The updated matrix consists of:
  • $\boldsymbol{K}^{\prime}$: a $k\times k$ ancestry relation matrix from the previous step,
  • $\mathit{rs}_{k}$: a $K\times 1$ vector of the indices of the $k$ splitting rules,
  • $\mathit{AR}_{i,j}$: the ancestry relation between rule where $r_{i}$ and $r_{j}$, where
    $\mathit{AR}_{i,j}=\begin{cases}1&\text{if }h_{i}\swarrow h_{j},\\ -1&\text{if }h_{i}\searrow h_{j},\\ 0&\text{if }h_{i}\overline{\left(\swarrow\vee\searrow\right)}h_{j}.\end{cases}$
The incremental nature of the $\mathit{calARM}$ is crucial, as it allows the ancestry matrix to be updated on-the-fly for each new rule. This property enables its integration directly into the sequential definition of $\mathit{nestedCombs}\left(K,G\right)$, avoiding inefficient post-hoc computation.
Prefix-closed filtering for nested combinations
Before incorporating $\mathit{update}_{\text{arMat}}$ into the definition of $\mathit{nestedCombs}\left(K,G\right)$, we first address the problem of eliminating configurations that contain a pair of crossed hyperplanes. According to Lemma (1), one approach is to check whether the symmetric elements with respect to the main diagonal of the ancestry relation matrix $\boldsymbol{K}$ contain a pair of zeros, i.e., $K_{i,j}=K_{j,i}=0$. However, this method only removes configurations after applying $\mathit{update}_{\text{arMat}}$, which wastes computational resources by updating configurations that would ultimately be invalid.
To resolve this, we introduce a prefix-closed filtering process over combinations. While conceptually similar to the prefix-closed filtering defined over tree datatypes in Part I, this version is tailored for nested combinations. It allows the post-hoc filtering to be fused directly into the incremental generation process of $\mathit{nestedCombs}\left(K,G\right)$, enabling the identification of crossed hyperplanes before partial configurations are extended to full ones. This provides an efficient solution for constructing a crossed-hyperplane-free nested combination generator.
Specifically, when the generator is defined sequentially, the following filter fusion theorem applies.
Theorem 2.
Filter fusion theorem. Let a sequential generator $\mathit{gen}$ be defined recursively as:
$\mathit{gen}$ $\left(\left[\>\right]\right)=\mathit{alg}_{1}\left(\left[\>\right]\right)$
$\mathit{gen}$ $\left(x:\mathit{xs}\right)=\mathit{alg}_{2}\left(x,\mathit{gen}\left(\mathit{xs}\right)\right),$
and consider a post-hoc filtering process. Then $\mathit{filtgen}_{q}=\mathit{filter}_{p}\circ\mathit{gen}$ can be fused into a single program defined as:
$\mathit{filtgen}_{q}$ $\left(\left[\>\right]\right)=\mathit{filter}_{q}\left(\mathit{alg}_{1}\left(\left[\>\right]\right)\right)$
$\mathit{filtgen}_{q}$ $\left(x:\mathit{xs}\right)=\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(x,\mathit{filtgen}_{q}\left(\mathit{xs}\right)\right)\right),$
provided that the fusion condition holds:
$\mathit{filter}_{p}\left(\mathit{alg}_{2}\left(x,\mathit{gen}\left(\mathit{xs}\right)\right)\right)=\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(x,\mathit{filter}_{p}\left(\mathit{gen}\left(\mathit{xs}\right)\right)\right)\right)$ (10)
In particular, if $\mathit{alg}_{2}:\mathcal{A}\times\left[\left[\mathcal{A}\right]\right]\to\left[\left[\mathcal{A}\right]\right]$ is defined as an extension operation that prepend $a:\mathcal{A}$ to $\mathit{as}:\left[\mathcal{A}\right]$ for all $\mathit{as}$ in $\mathit{ass}:\left[\left[\mathcal{A}\right]\right]$, then proving the fusion condition (10) is equivalent to proving the prefix-closed property:
$p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$ (11)
Proof.
We prove the fusion theorem by following reasoning
$\mathit{filtgen}_{q}$
$=$ $\mathit{filter}_{p}\circ\mathit{gen}$
$=$ $\mathit{filter}_{p}\left(\mathit{alg}_{2}\left(x,\mathit{gen}\left(\mathit{xs}\right)\right)\right)$
[…]
$\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(x,\mathit{filter}_{p}\left(\mathit{gen}\left(\mathit{xs}\right)\right)\right)\right)$
$=$ $\{\text{definition of $\mathit{filtgen}_{q}$}\}$
$\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(x,\mathit{filtgen}_{q}\left(\mathit{xs}\right)\right)\right)$
The equivalence between filter fusion condition (10) and prefix-closed property (11) is straight forwards. A configuration $a:\mathit{as}$ survived in $\mathit{filter}_{p}$ will also survived in $\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(a,\mathit{filter}_{p}\left(\left[\mathit{as}\right]\right)\right)\right)$ because $p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$. ∎
The reason the prefix-closed property (11) introduces an additional predicate $q$ is that, if we already know $p\left(\mathit{as}\right)$ holds, it is often more efficient to evaluate $q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$ rather than directly computing $p\left(a:\mathit{as}\right)$. This explains why the fused generator $\mathit{filtgen}_{q}$is more efficient than the post-hoc approach $\mathit{filter}_{p}\circ\mathit{gen}$. For example, in the classical eight queens problem, $p$ checks that no queen attacks any other, while the auxiliary predicate $q$ only verifies that the newly added queen does not attack the others.
Similarly, for our problem, since $\mathit{nestedCombs}\left(K,D\right)$ can be defined sequentially, we can establish a prefix-closed property for the ancestry relation matrix, analogous to the auxiliary check in the eight queens problem.
Fact 2.
Prefix-closed property for crossed hyperplanes. Assume we have a feasible combination of hyperplanes $\mathit{hs}_{K}$ (i.e., $p\left(\mathit{hs}\right)=\text{True}$), when adding a new hyperplane $h$ to $\mathit{hs}_{K}$ the following prefix-closed property holds:
$p\left(h:\mathit{hs}_{K}\right)=q\left(h:\mathit{hs}_{K}\right)\wedge p\left(\mathit{hs}_{K}\right),$
where
$q\left(h:\mathit{hs}_{K}\right)=\lnot\exists h_{i}\in\mathit{hs}_{K}:p_{\text{crs}}\left(h,h_{i}\right)=\text{True},$
i.e., $q\left(h:\mathit{hs}_{K}\right)=\text{True}$ if $h$ does not cross any hyperplane.
The naive predicate $p\left(h:\mathit{hs}_{K}\right)$ requires checking all pairs of hyperplanes in $h:\mathit{hs}_{K}$ for crossings, which has complexity $O\left(\left(K+1\right)^{2}\times t\right)$, where $t$ is the cost of evaluating $p_{\text{crs}}$. Fact 2 shows that, since $\mathit{hs}_{K}$ is already feasible ($p\left(\mathit{hs}_{K}\right)=\text{True}$), it suffices to check only whether the newly added hyperplane $h$ crosses any existing $h_{i}\in\mathit{hs}_{K}$. This reduces the complexity to $O\left(K\times t\right)$.
  1. 1.
    Input: $r_{j}$: the index of the new splitting rule; $\mathit{ncsr}$: a $M\times\left(k+1\right)\times k$ tensor, which contains a vector of $\mathit{ncr}:NCR$ configurations; $\mathit{css}$: a matrix for storing all $G$-combinations of data points of all splitting rules; $\mathit{asgn}^{\pm}$: the positive/negative predictions of each splitting rule
  2. 2.
    Output: The updated nested combinations $\mathit{ncsr}^{\prime}:\mathit{NCRs}$ without crossed hyperplanes, which is a a vector of $\left(k+2\right)\times\left(k+1\right)$ matrices
  3. 3.
    $\mathit{ncrs}^{\prime}=\left[\>\right]$
  4. 4.
    for $\mathit{ncr}\in\mathit{ncrs}$ do:
  5. 5.
    for $r_{i}\in\mathit{ncr}\left[0\right]$ do:
  6. 6.
       $p_{1}=\text{True }\text{if }\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{+}\left[r_{i}\right]\vee\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{-}\left[r_{i}\right]$ else False // the $G$-combination of data points $\mathit{css}\left[G\right]\left[r_{j}\right]$ all lies in the positive side or negative side of the hyperplane $r_{i}$,
  7. 7.
      $p_{2}=\text{True }\text{if }$ $\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{+}\left[r_{j}\right]\vee\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{-}\left[r_{j}\right]$ else False // the $G$-combination corresponds for defining rule $r_{i}$ all lies in the positive side or negative side of the splitting rule $r_{j}$, $\mathit{unrank}\left(r_{i}\right)$ takes an index and return all data points for defining splitting rule $r_{i}$,
  8. 8.
      if $p_{1}\vee p_{2}$ do:
  9. 9.
       $\mathit{ncr}^{\prime}=\mathit{update}_{\text{arMat}}\left(\mathit{ncr}\right)$ // update ancestry relation matrix
  10. 10.
       $\mathit{ncrs}^{\prime}+\!\!+\left[\mathit{ncr}^{\prime}\right]$
  11. 11.
      else:
  12. 12.
        break // break the inner loop because $\mathit{ncr}$ is infeasible
  13. 13.
    return $\mathit{ncrs}^{\prime}$
Algorithm 1 $\mathit{updates}_{\text{arMat}}$
We can now incorporate $\mathit{update}_{\text{arMat}}$ along with the additional filtering process for eliminating crossed hyperplanes directly into the definition of $\mathit{nestedCombs}\left(K,G\right)$. For simplicity, we define a batched version of $\mathit{update}_{\text{arMat}}$, denoted as $\mathit{updates}_{\text{arMat}}$, that operates on a list of configurations $\mathit{ncrs}$, as implemented in Algorithm 1.
With these components, we can construct the incremental, crossed-hyperplane-free ancestry relation matrix generator $\mathit{\mathit{nestedCombsFA}}$. This is achieved by introducing a single additional line of code after line 12 of Algorithm using Algorithm 1 to update all feasible nested combinations while avoiding non-feasible combinations containing crossed hyperplanes. The resulting modification is minimal—just one line—and the complete pseudocode is provided in Appendix A.2, Algorithm 6.
3.3 Optimal hypersurface decision tree algorithm
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: number of splitting rules; $G$: dimension of the embedded space
  2. 2.
    Output: Array of $\left(k,D\right)$-nested-combinations for $k\in\left\{0,\ldots K\right\}$
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    $\rho_{M}\left(\mathit{xs}\right)=\mathit{embed}\left(M,\mathit{xs}\right)$ // map the data list $\mathit{xs}$ to the embedded space
  7. 7.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  8. 8.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  9. 9.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  10. 10.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  11. 11.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  12. 12.
    $\mathit{css}\left[G\right]=\left[\>\right]$
  13. 13.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  14. 14.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  15. 15.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  16. 16.
        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$
  17. 17.
[…]
  18. 18.
    if $E_{\text{0-1}}\left(t_{\text{best}}\right)\leq E_{\text{0-1}}\left(t_{\text{opt}}\right)$:
  19. 19.
      $t_{\text{opt}}=t_{\text{best}}$
  20. 20.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$ // eliminate size $K$ nested combinations after evaluation
  21. 21.
    return $t_{\text{opt}}$
Algorithm 2 $\mathit{hodt}\left(K,M,\mathit{xs}\right)$
Finally, after introducing the prefix-closed filtering and incremental update for ancestry relation matrix, we are now ready to fuse all components into a single program. In particular, the specification 5 can be solved exactly using the program implemented in Algorithm 3.
The main difference between the brute-force specification 5 and Algorithm 3 lies in the order of execution. In the specification, functions are composed sequentially, meaning each function begins only after the previous one has completed. If written as pseudocode, both $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ would appear outside the outermost loop that defines $\mathit{\mathit{nestedCombs}}\left(K,G\right)$. This is inefficient because $\mathit{\mathit{nestedCombs}}\left(K,G\right)$ generates a prohibitively large number of candidate solutions.
By fusing the composition $\mathit{min}{}_{E_{\text{0-1}}}\circ\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}\circ\mathit{\mathit{nestedCombsFA}}$ into a single program—placing $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ inside the outermost loop, e.g., “for $n\leftarrow\mathit{range}\left(0,N\right)$ do,”—we can solve the problem in a single recursive pass. Since we are designing an optimal algorithm, every seemingly simple modification to the brute-force program must be rigorously justified; otherwise, optimality could be lost. Thus, the seemingly straightforward solution in Algorithm 3 does not rely on low-level implementation tricks but stems from a deep understanding of the algorithmic structure of the problem.
Unlike ad-hoc methods such as branch-and-bound algorithms, we formalize the problem as a brute-force program and then derive a correct-by-construction implementation. Achieving this requires careful definition and integration of various generators—decision tree, splitting rule, ancestry relation generators—along with proofs of correctness for each fusion step: from reducing the original problem to a simplified decision tree problem, to dynamic programming fusion, nested combination fusion, ancestry relation matrix construction, and prefix-closed filtering fusion. Although formal reasoning adds complexity to algorithm design, it is essential for the development of exact algorithms.
We hope that the rigor invested in designing this optimal algorithm will positively influence future studies, demonstrating not only how formal steps can guarantee optimality but also the advantages of adopting a formal, principled approach.
4 Heuerestic methods
The use of heuristics is common in the study of optimal algorithms. Examples include setting a time limit with random initialization (a variant developed by Dunn (2018)), employing depth-first search with a time limit [^1] Although some BnB algorithms claim that using depth-first search with a time limit does not prevent finding the optimal solution if the algorithm continues running, this claim is vacuous in practice: without a feasible running-time bound, finding the optimal solution may still require exponential time in the worst case. (Hu et al., 2019; Lin et al., 2020), or using binarization for continuous data Brița et al. (2025). It should be noted that the purpose of these heuristics is not to demonstrate superiority—unless accompanied by a rigorous performance guarantee, which would constitute a different line of research—but rather to provide quick, plausible solutions. Each heuristic has its own strengths and weaknesses, but the shared goal is to obtain a plausible solution quickly when the combinatorial complexity of the problem is astronomical, making exact solutions computationally or memory-wise intractable.
For example, a depth-first search with a time limit is particularly effective when the tree size is large: it can quickly produce a plausible solution due to the flexibility afforded by a large tree. In contrast, for problems requiring a strictly small tree (e.g., 3–4 leaves), the depth-first search strategy is less effective, as it explores solutions more slowly than a breadth-first approach and is more difficult to parallelize.
Similarly, the HODT problem exhibits formidable combinatorics. Even for a modest dataset with $N=100$, $D=3$, $K=3$, there are $\left(\begin{array}[]{c}\left(\begin{array}[]{c}100\\ 3\end{array}\right)\\ 3\end{array}\right)\times 3!\approx 4\times 10^{15}$ possible decision trees in the worst case. Solving this problem exactly is currently intractable for our algorithms. In this section, we develop two simple yet effective heuristics for addressing the HODT problem.
  1. 1.
    Input: Parameteres for coreset: $\mathit{BS}$: Block size; $R$: number of shuffle time in each filtering process; $L$: Max-heap size; $B_{\max}$: Maximum input size for the Deep-ICE algorithm; $c\in\left(0,1\right]$: Shrinking factor for heap size
  2. 2.
      Parameteres for $\mathit{hodt}$: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\mathit{xs}$: input data list;
  3. 3.
    Output: Max-heap containing top $L$ configurations and associated data blocks
  4. 4.
    Initialize coreset $\mathcal{C}\leftarrow\mathit{ds}$
  5. 5.
    while $\mathcal{C}\leq B_{\max}$ do:
  6. 6.
    Reshuffle the data, divide $\mathcal{C}$ into $\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil$ blocks $\mathcal{C}_{B}=\left\{C_{1},C_{2},\ldots,C_{\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil}\right\}$
  7. 7.
    Initialize a size $L$ max-heap $\mathcal{H}_{L}$
  8. 8.
    for $r\leftarrow 1$ to $R$ do:
  9. 9.
      $r=r+1$
  10. 10.
      for $C\in\mathcal{C}_{B}$ do:
  11. 11.
       $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,C\right)$
  12. 12.
       $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},C\right)$
  13. 13.
    $\mathcal{C}\leftarrow\mathit{unique}\left(\mathcal{H}_{L}\right)$ // Merge blocks and remove duplicates
  14. 14.
    $L\leftarrow L\times c$ // Shrink heap size:
  15. 15.
    $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,\mathcal{C}\right)$ // Final refinement
  16. 16.
    $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},\mathcal{C}\right)$
  17. 17.
    return $\mathcal{H}_{L}$
Algorithm 3 $\mathit{hodt}Coreset\left(K,M,\mathit{xs},\mathit{BS},R,L,B_{\max},c\right)$
4.1 Coreset selection method
The first method, as reported by He et al. (2025), has demonstrated superior performance for empirical risk minimization in two-layer neural networks. The method is based on the following idea: instead of computing the exact solution across the entire dataset—which is computationally infeasible for large $K$ and DD—the coreset method identifies the exact solution for the most representative subsets.
The coreset method functions as a “layer-by-layer” data filtering process, in each loop, we ran the $\mathit{hodt}$ algorithm for a subset of the dataset, and only keep the $L$ best solutions with respect to the whole datasets. Since better configurations tend to have lower training accuracy, they are more likely to “survive” during the selection process. By recursively reduces the data size until the remaining subset can be processed by running the complete optimal algorithm. The algorithm process is detailed in Algorithm 3.
4.2 $\mathit{sodt}$ with selected hyperplanes
  1. 1.
    Input: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\alpha$: threshold for duplicate data $\mathit{xs}$: input data list;
  2. 2.
    Output: The updated nested combinations $\mathit{NCs}$ without crossed hyperplanes, a vector of e $\left(k+2\right)\times\left(k+1\right)$ matrices
  3. 3.
    $t_{\text{best}}=\mathit{empty}\left(K,1\right)$ // a vector
  4. 4.
    $\mathcal{H}_{C}=\mathit{\mathit{hodt}Coreset}\left(1,M,\mathit{xs},BS,R,L,B_{\max},c\right)$ // selecting candidate hypersurfaces with low 0-1 loss
  5. 5.
    $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$
  6. 6.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  7. 7.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  8. 8.
    for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,n+1\right)\right)$ do:
  9. 9.
      $\mathit{ncss}^{\prime}=\mathit{filter}_{q\left(\alpha\right)}\left(\mathit{map}\left(\cup\left[n\right],\mathit{ncss}\left[k-1\right]\right)\right)$ // filter out nested combinations $\mathit{nc}$ in $\mathit{ncss}^{\prime}\left[k\right]$ with duplicates data points small than threshold $\alpha$
  10. 10.
      $t_{\text{best}}\left[k\right]=\mathit{min}_{E_{\text{0-1}}}\left(\mathit{mapL}\left(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}^{\prime}\right)\right)$
  11. 11.
      $\mathit{ncss}\left[k\right]=\mathit{ncss}\left[k\right]\cup\mathit{reverse}\left(\mathit{ncss}^{\prime}\right)$ // update $\mathit{ncss}$ to generate combinations in revolving door ordering,
  12. 12.
    $\mathit{ncss}\left[K\right]=\left[\>\right]$
  13. 13.
    return $t_{\text{best}}$
Algorithm 4 $\mathit{sodtWSH}\left(K,M,\alpha,\mathit{xs}\right)$
Although the coreset selection method provides plausible solutions for low-dimensional datasets, an obvious limitation arises for high-dimensional datasets: the number of possible splitting rules is too large, even when the dataset is partitioned into small blocks. Moreover, defining a decision tree with $K$ splitting rules requires at least $K\times D$ distinct data points. Running $\mathit{hodt}$ for $K\times D$ points becomes intractable for most high-dimensional datasets. For example, when $K=2$, and $D=20$, the block size is $2\times 20$ and the number of possible hyperplanes is $\left(\begin{array}[]{c}40\\ 25\end{array}\right)\approx 1\times 10^{11}$, which is prohibitively large. In practice, our current implementation can efficiently process at most $D+2$ or $D+3$ points when $D\geq 20$, far below the ideal block size of $K\times D$.
To address this limitation, we leverage a key advantage of the size-constrained $\mathit{hodt}$ algorithm—its ability to decompose the original difficult ODT problem into many smaller subproblems, each of which can be solved efficiently using $\mathit{sodt}$. Solving the full ODT problem directly is computationally prohibitive due to the enormous combinatorial space. Instead, we focus on finding a set of “good candidate” hyperplanes, and then construct decision trees from these candidates rather than exploring the entire search space.
This insight motivated the development of a new heuristic, $\mathit{sodtWSH}$, (short for $\mathit{sodt}$ with selected hyperplanes”), described in Algorithm 4. The algorithm identifies candidate hyperplanes with relatively low training loss and constructs optimal decision trees based on these hyperplanes. Rather than generating hyperplanes sequentially from continuous data blocks—where each hyperplane differs from the previous by only one or two points—we generate a large set of candidate hyperplanes and then apply $\mathit{sodt}$ only to combinations in which the data points defining each splitting rule are sufficiently distinct—we want to only evaluate those that contains unique data points greater than a threshold $\alpha<D\times K$. This will helps us to find sufficiently distinct hyperplanes.
5 Experiments
The experiments aim to provide a detailed analysis along four dimensions:
  1. 1.
    Computational complexity and scalability: We compare $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{rec}}$ under both sequential and parallelized settings, evaluating the scalability of the more efficient method.
  2. 2.
    Empirical combinatorial complexity: We analyze the combinatorial complexity of hyperplane and hypersurface decision tree models, demonstrating that the true complexity—after filtering out crossed hyperplanes—is substantially smaller than the theoretical upper bound provided in Part I.
  3. 3.
    Analysis over synthetic datasets: Using synthetic datasets generated from hyperplane decision trees (with data lying in convex polygon regions), we benchmark the performance of $\mathit{\mathit{hodt}Coreset}$. We systematically test the effects of ground truth tree size, data dimensionality, dataset size, label noise, and data noise separately. Our results show that the hyperplane decision tree model learned by our algorithm is not only more accurate in prediction but also more robust to noise.
  4. 4.
    Generalization performance on real-world datasets: We evaluate performance across 30 real-world datasets, comparing our hyperplane decision tree model learned by $\mathit{sodtWSH}$ against the state-of-the-art optimal decision tree algorithm Brița et al. (2025) and the well-known approximate algorithm CART. We demonstrate that, when model complexity is properly controlled, hyperplane decision trees consistently outperform axis-parallel models in out-of-sample tests.
As discussed, the combinatorial complexity of the HODT problem is currently intractable even for moderately sized datasets. Therefore, the experiments for points (3) and (4) are conducted exclusively for the $M=1$ (hyperplane) case. To obtain high-quality solutions efficiently, we employ the two heuristic methods developed in Section 4.
In these experiments, we adapt Brița et al. (2025)’s ConTree algorithm rather than Mazumder et al. (2022)’s Quant-BnB algorithm. We observed that ConTree often provides more accurate solutions while being considerably more efficient, making it more suitable for our comparisons. We note that the current experiments do not explore the effects of hyperparameter tuning, such as minimum data per leaf or tree depth when optimizing size-constrained decision trees. The performance impact of these fine-grained controls represents an interesting avenue for future research.
All experiments were conducted on an Intel Core i9 CPU with 24 cores (2.4–6 GHz), 32 GB RAM, and a GeForce RTX 4060 Ti GPU.
5.1 Computational analysis
In this section, we analyze in detail the computational efficiency of $\mathit{sodt}$ for solving the size-constrained ODT problem. Since $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ share similar advantages—both are fully vectorizable but unable to exploit efficient dynamic programming (DP)—we focus on comparing $\mathit{sodt}_{\text{vec}}$ against the dynamic programming algorithm $\mathit{sodt}_{\text{rec}}$.
Rather than using a binary tree data structure, which relies on pointers to locate subtrees and suffers from poor cache performance, both $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{rec}}$ are implemented entirely with array (heap) data structures. This design ensures contiguous memory allocation, significantly reducing cache misses.
We compare their performance in two settings: 1) Sequential setting: Nested combinations are processed one-by-one using a standard for-loop. 2) Parallelized setting. A large batch of feasible nested combinations is processed simultaneously. Specifically, we store $\mathit{ncrs}:\mathit{NCRs}$ in a single large tensor instead of a list and pass it as input to $\mathit{mapL}\left(\mathit{sodt},\mathit{ncrs}\right)$. For $\mathit{mapL}\left(\mathit{sodt}_{\text{vec}}\right)$, we implement this as a single function (batch_sodt). This allows us to efficiently process the large batch of feasible nested combinations $\mathit{ncrs}$ on both CPU and GPU. We denote the results as $\mathit{sodt}_{\text{vec}}^{\text{cpu}}$ and $\mathit{sodt}_{\text{vec}}^{\text{gpu}}$, respectively.
In contrast, since $\mathit{sodt}_{\text{rec}}$ cannot be fully vectorized, we parallelize $\mathit{mapL}\left(\mathit{sodt}_{\text{vec}}\right)$ using multi-core CPU execution, initialized in a multi-process setting.
5.1.1 Comparison between the vectorized and recursive implementation
Figure 5: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on sequential setting. .
Figure 6: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on parallel setting..
Even though $\mathit{sodt}_{\text{rec}}$ has provably lower theoretical complexity, the hardware compatibility of $\mathit{sodt}_{\text{vec}}$ allows it to outperform in practice, particularly as $K$ increases. As shown in Figure 5, in the sequential setting $\mathit{sodt}_{\text{rec}}$ is the most efficient method when $K\leq 3$. However, it becomes slower than $\mathit{sodt}_{\text{vec}}^{\text{cpu}}$ at $K=4$, and eventually the slowest method once $K\geqq 5$. In contrast, $\mathit{sodt}_{\text{vec}}^{\text{gpu}}$ is initially slower than all other methods but overtakes $\mathit{sodt}_{\text{rec}}$ after $K=5$, ultimately becoming the most efficient method at $K=7$.
In the parallel setting (Figure 6), the benefits of vectorization are even more pronounced. Parallelizing $\mathit{mapL}\left(\mathit{sodt}_{\text{rec}}\right)$ with multiprocessing incurs significant overhead from initializing multiple CPU cores, making $\mathit{sodt}_{\text{rec}}$ consistently the slowest method across all cases. Comparing $\mathit{sodt}_{\text{vec}}^{\text{cpu}}$ and $\mathit{sodt}_{\text{vec}}^{\text{gpu}}$, we observe that when $K<4$, the CPU implementation is more efficient. For $K=4$, both implementations achieve nearly identical performance, while for $K>4$, $\mathit{sodt}_{\text{vec}}^{\text{gpu}}$ becomes superior. The GPU computation does not always outperform CPU computation because $\mathit{mapL}\left(\mathit{sodt}_{\text{vec}}\right)$ in its current implementation consumes excessive memory to store intermediate results. The resulting memory transfer overhead can outweigh computational gains when $k$ is small. A more sophisticated low-level implementation (e.g., in CUDA rather than Python) may reduce these overheads and further improve GPU performance.
5.1.2 Computational scalability of the vectorized method—the ability to explore one million nested combinations within a fixed time
Figure 7: Log-log wall-clock run time (seconds) for the $\mathit{sodt}_{\text{vec}}$ algorithm, across nested combinations of size up to $1\times 10^{6}$. On this scale, linear run time appears as a logarithmic function of problem size.
Following the detailed comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ in both sequential and parallel settings, we now turn to the computational scalability of the winning method, $\mathit{sodt}_{\text{vec}}$. We focus on analyzing the pure computational scalability of the algorithm by benchmarking its performance without any acceleration techniques. Specifically, we measure the total wall-clock runtime of $\mathit{sodt}_{\text{vec}}$ for solving $1\times 10^{6}$ instances of feasible nested combinations (with no crossed hyperplanes) for different values of KK。In our setup, cases with $K\leq 4$ are executed on the CPU, while cases with $K>4$ are executed on the GPU.
As expected, when $K$ is fixed $\mathit{sodt}_{\text{vec}}$ (with worst-case complexity $O\left(K!\times N\right)$) exhibits linear runtime growth, which appears logarithmic in a log-linear plot. The batched vectorized implementation proves highly efficient for $K=2,3$; for example, it solves $1\times 10^{6}$ feasible nested combinations in only a few tens of seconds. Note that a single size $K$ nested combination can generate up to $K!$ possible proper decision trees in the worst-case. These results demonstrate the clear computational advantages of the $\mathit{sodt}_{\text{vec}}$ algorithm.
All current experiments are implemented in Python using the PyTorch library, and we anticipate that a lower-level implementation will further enhance performance.
5.2 Combinatorial complexity of the hyperplane/hypersurface decision tree
5.2.1 Combinatorial complexity of nested combinations after fitlered out crossed hyperplanes
Figure 8: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
As noted, the combinatorial complexity of the hyperplane ($M=1$) ODT problem is bounded by $\left(\begin{array}[]{c}\left(\begin{array}[]{c}N\\ D\end{array}\right)\\ K\end{array}\right)=O\left(N^{DK}\right)$. However, as established in Theorem 1, any nested combination containing a pair of crossed hyperplanes cannot yield a complete decision tree. Consequently, all such infeasible nested combinations can be filtered out without compromising the optimality of the algorithms.
To evaluate the true combinatorial complexity of the problem after excluding infeasible combinations, we conducted experiments on synthetic Gaussian datasets of size $N=2000$. Due to memory limitations, we tested only subsets of the original dataset (denoted as “blocksize” in Figure 8. Specifically, a small sub-dataset was sampled from the $N=2000$ dataset and then passed into $\mathit{nestedCombsFA}$, while hyperplane predictions were still carried out using the full $N=2000$ dataset, but blocksize ranged from 8 to 10.
For each panel in Figure 8, we fixed the dimension $D$ and varying $N$ and $K$. We computed the combinatorial complexity over five datasets and reported the mean and standard deviation, illustrated as bar charts with error bars. The results show that the true combinatorial complexity (bars), after filtering infeasible nested combinations, is substantially smaller than the theoretical upper bound $O\left(N^{DK}\right)$ (black lines).For example, when $\text{blocksize}=10$, $D=2$, $K=10$, the theoretical number of nested combinations is $3\times 10^{9}$, whereas the number of feasible nested combinations is only $1\times 10^{7}$, more than 300 times difference!
An interesting pattern emerges in the $D=2$ panel: the complexity curve forms an inverted U-shape, indicating that the combinatorial complexity of decision trees initially increases with $K$ but eventually decreases. This phenomenon aligns with the explanation given in He and Little (2025), under mild probabilistic assumptions, the likelihood of constructing a feasible nested combination decreases exponentially with $K$. The inverted U-shape arises because, at first, the rapid growth in the number of nested combinations dominates, but beyond a certain threshold, the exponential decay in feasibility prevails. Nevertheless, in practice, this behavior is rarely observed, as it typically requires $K$ to reach astronomically large values even for medium-sized datasets.
Comparing the combinatorial complexity of hypersurface and hyperplanes in the same diemension
Figure 9: Combinatorial complexity of feasible nested combinations with varying $K$. Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
Having analyzed the combinatorial complexity of hyperplane splitting rules, we now extend the analysis to hypersurface rules by comparing their complexity with that of hyperplanes in the same dimension. As discussed, a hypersurface defined by a degree-$M$ polynomial can be represented as a hyperplane embedded in a space of dimension $G=\left(\begin{array}[]{c}D+M\\ D\end{array}\right)-1$ dimensional space.
For instance, when $D=4$ and $M=2$ we obtain $G=14$. Our goal is to determine whether hypersurfaces in $\mathbb{R}^{D}$ (corresponding to hyperplanes in $\mathbb{R}^{G}$) exhibit different combinatorial complexity compared to hyperplanes. As illustrated in Figure 9, the degree-$M$ hypersurface in $\mathbb{R}^{D}$ shows slightly higher combinatorial complexity than the corresponding hyperplane in $\mathbb{R}^{G}$, This increase may stem from the greater expressivity of hypersurfaces relative to hyperplanes.
5.3 Computational experiments with synthetic datasets
In this section, we evaluate the performance of the HODT model on a variety of synthetically generated datasets. The aim is to assess how effectively a more flexible model can recover the underlying ground truth, which is itself generated by a decision tree model.
Our experiments build on those of Murthy and Salzberg (1995) and Bertsimas and Dunn (2017). While their datasets were generated using axis-parallel decision trees, we generalize the setting by using hyperplane decision trees. Consequently, the underlying partitions in our datasets are polygonal regions. Specifically, we generate synthetic datasets from randomly constructed hyperplane decision trees, and then compare the performance of different algorithms in inducing trees that approximate the ground truth.
To construct the ground truth, we generate a decision tree of a specified size by choosing splits at random. The leaves of the tree are assigned unique labels so that no two leaves share the same label, ensuring the tree is the minimal representation of the ground truth. Training and test datasets are then generated by sampling each data point $x$ uniformly at random and assigning its label according to the ground truth tree. In each experiment, 50 random trees are generated as ground truths, producing 50 corresponding training-test dataset pairs. The training set size varies across experiments, while the test set size is fixed a $2^{d}\times\left(D-1\right)\times 500$, where $d$ is the depth of the tree and $D$ s the data dimensionality.
Following Murthy and Salzberg (1995); Bertsimas and Dunn (2017), we evaluate tree quality using six metrics:
  • Training accuracy: Accuracy on the training set.
  • Test accuracy: Accuracy on the test set.
  • Tree size: Number of branch nodes.
  • Maximum depth: Maximum depth of the leaf nodes.
  • Average depth: Mean depth of leaf nodes.
  • Expected depth: Average depth of leaf nodes weighted by the proportion of the test set samples assigned to each leaf.
We also follow Bertsimas and Dunn (2017) in examining the effect of five sources of variation or noise: 1) Ground-truth tree size, 2)Training data size, 3)Label noise, 4) Feature noise, 5) Data dimensionality.
Since Bertsimas and Dunn (2017)’s implementation is not publicly available, and replicating their construction is difficult, so we benchmark against three widely used axis-parallel decision tree algorithms: CART-depth (depth-constrained CART algorithm), CART-size (size-constrained CART algorithm), and the state-of-the-art optimal axis-parallel ODT algorithm—ConTree (Brița et al., 2025) (which only allows tree depth to be specified). For all experiments in this section, we employ the $\mathit{hodtCoreset}$ algorithm for generating accurate hypersurface decision tree.
Effect of tree size
Tree size ($K$) Method Train Acc (%) Test Acc (%) Tree size Depth
Maximum Average Expected
2
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.12 (0.04)
93.82 (3.01)
91.72 (3.83)
100 (0.00)
100 (0.00)
82.61 (0.04)
86.12 (3.73)
84.57 (4.03)
93.50 (3.21)
100 (0.00)
2.92 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.49 (0.16)
2.00 (0.00)
1.67 (1.20)
1.67 (1.20)
1.96 (0.15)
2.15 (0.20)
2.00 (0.00)
1.65 (0.16)
1.66 (0.16)
3
CART-depth-3
CART-size-6
ConTree-3
HODT
Ground Truth
90.72 (047)
93.88 (3.89)
95.98 (2.93)
100 (0.00)
100 (0.00)
81.30 (5.71)
83.92 (5.55)
85.25 (4.96)
94.91 (2.36)
1 (0.00)
5.90 (0.00)
6.00 (0.00)
6.94 (0.24)
3.00 (0.00)
3.00 (0.00)
3.00 (0.00)
4.02 (0.65)
3.00 (0.00)
3.00 (0.00)
2.72 (0.45)
2.83 (0.14)
3.10 (0.24)
2.99 (0.03)
2.25 (0.0)
2.18 (0.11)
2.77 (0.19)
2.71 (0.20)
2.99 (0.04)
2.18 (0.25)
2.07 (0.24)
4
CART-depth-3
CART-size-8
ConTree-3
HODT
Ground Truth
87.98 (5.26)
93.90 (3.93)
93.12 (4.02)
97.46 (1.70)
100 (0.00)
77.66 (5.38)
81.93 (4.93)
81.51 (4.78)
89.89 (3.71)
100 (0.00)
6.12 (90.86)
8.00 (0.00)
7.00 (0.00)
4.00 (0.00)
4.00 (0.00)
3.00 (0.00)
4.54 (0.57)
3.00 (0.00)
3.52 (0.50)
3.52 (0.50)
2.87 (0.14)
3.49 (0.21)
3.00 (0.00)
2.63 (0.19)
2.63 (0.19)
2.86 (0.15)
3.13 (0.21)
3.00 (0.00)
2.60 (0.31)
2.63 (0.33)
Table 1: The effect of tree size of ground truth tree. No noise in data, training size = 100.
In our first experiments, we evaluated the effectiveness of each method as problem complexity increased. We fixed the training set size at $N=100$ and $D=2$, while varying the ground truth size $K$ from 2 to 4. Table 1 presents the results. Since our goal was to solve the size-constrained ODT problem, we set the depth of the depth-constrained algorithms as $d=\left\lceil\log_{2}\left(K\times D\right)+1\right\rceil$ for the CART-depth and ConTree algorithms. CART-size constrained the tree size to $K\times D$. Note that a depth $d$ tree has at most $2^{d}-1$ tree size.
The results show that our algorithm has a very high probability of finding the optimal solution (100% for $K=2$ and $K=3$) and HODT significantly outperforms the other methods. Although all methods show a decrease in both training and test accuracy as $K$ increase to 4, this decline is expected due to the increasing problem complexity. Even so, HODT and ConTree are more robust to changes in tree size. When $K$ increases to 4, CART-size and CART-depth exhibit a larger reduction in out-of-sample accuracy (4.19% and 4.95%, respectively) compared to ConTree and HODT, which show smaller decreases (3.06% and 3.61%). Interestingly, although the optimal axis-parallel decision tree obtained by the ConTree algorithm is more accurate than CART-depth under the same depth constraint, ConTree often fully exploits the given depth, resulting in nearly maximal tree sizes in most cases. Consequently, ConTree typically produces larger trees than CART-depth. However, CART-size, by allowing just one additional leaf, achieves better solutions than ConTree by providing greater flexibility in tree depth.
In summary, the experiments demonstrate that HODT is not only more accurate than CART and ConTree across varying tree sizes and depths, but also closely matches the ground truth across all measures. In contrast, axis-parallel trees learned using CART and ConTree are significantly larger than HODT trees and produce substantially worse solutions.
Effect of data size
Training set size Method Train Acc (%) Test Acc (%) Tree size Depth
Maximum Average Expected
100
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.76 (5.11)
97.62 (2.51)
93.52 (3.53)
100 (0.00)
100 (0.00)
75.46 (15.70)
80.26 (16.80)
78.62 (16.14)
84.38 (19.27)
100 (0.00)
2.88 (0.32)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
0.00 (0.00)
2.00 (0.00)
3.92 (0.69)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.96 (0.11)
3.01 (0.30)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.96 (0.12)
2.52 (0.33)
2.00 (0.00)
1.66 (0.25)
1.70 (0.25)
200
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.00 (4.49)
94.49 (3.02)
92.87 (3.38)
99.94 (0.16)
100 (0.00)
84.62 (5.23)
89.57 (4.04)
88.51 (3.99)
97.97 (1.01)
100 (0.00)
2.92.00 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.50 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.98 (0.07)
2.17 (0.21)
2.00 (0.00)
1.71 (0.25)
1.70 (0.25)
400
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
87.45 (4.69)
92.58 (3.48)
91.44 (3.37)
99.85 (0.21)
100 (0.00)
84.62 (4.90)
89.35 (4.08)
88.78 (4.00)
98.98 (0.58)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.53 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.05)
2.15 (0.24)
2.00 (0.00)
1.71 (0.25)
1.70 (0.25)
800
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
86.91 (4.63)
92.21 (3.43)
91.00 (3.59)
99.87 (0.14)
100 (0.00)
85.14 (4.96)
90.12 (3.82)
89.18 (3.94)
99.40 (0.36)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.3 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.53 (0.18)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.13 (0.22)
2.00 (0.00)
1.69 (0.25)
1.70 (0.25)
1600
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
85.66 (5.05)
91.48 (3.69)
90.13 (3.69)
99.87 (0.12)
100 (0.00)
84.97 (5.15)
90.53 (3.71)
89.38 (3.84)
99.65 (0.22)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.51 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.16 (0.26)
2.00 (0.00)
1.70 (0.24)
1.70 (0.25)
Table 2: The effect of training data size. Ground truth-tree size $K=2$.
The second set of experiments demonstrates the effect of training data size relative to problem complexity. We increased the size of the training set while keeping the ground truth size fixed at $K=2$ and the dimension at $D=2$. ConTree and CART-depth were restricted to depth 2, while CART-size was restricted to $K\times D=4$.
No noise was added to the data. Table 2 shows that out-of-sample accuracy increased for all methods as the training set grew. The improvement was most pronounced for HODT, which achieved a 15.27% increase in out-of-sample accuracy, whereas ConTree, CART-size, and CART-depth improved by 10.76%, 10.27%, and 9.51%, respectively. Notably, CART-depth performed significantly worse than the other methods.
These results demonstrate that even in data-poor environments, optimizing an appropriate model (HODT in this case) substantially improves out-of-sample performance. Even when axis-parallel trees do not match the ground truth, optimizing the solution to optimality still produces significant differences. The optimal algorithm (ConTree) achieves much higher test accuracy than the approximate method (CART-depth) on both training and test datasets. This observation is consistent with Bertsimas and Dunn (2017), and our experiments provide additional evidence for scenarios in which the ground truth does not align with the chosen model. These results offer clear support against the notion that optimal methods necessarily overfit the training data in data-scarce settings.
Effect of data dimension
Data dimension Method Train Acc (%) Test Acc (%) Tree size Depth
Maximum Average Expected
2
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.00 (4.49)
94.49 (3.02)
92.87 (3.38)
99.94 (0.16)
100 (0.00)
84.62 (5.23)
89.57 (4.04)
88.51 (3.99)
97.97 (1.01)
100 (0.00)
2.92 (0.27)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.53 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.98 (0.07)
2.17 (0.21)
2.0 (0.00)
1.71 (0.25)
1.70 (0.25)
4
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
79.94 (4.53)
84.17 (4.41)
82.80 (4.16)
95.13 (1.38)
100 (0.00)
73.97 (4.74)
76.72 (4.48)
76.16 (4.82)
90.97 (2.55)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.10 (0.3)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.48 (0.13)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.24 (0.18)
2.00 (0.00)
1.67 (0.13)
1.67 (0.12)
8
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
75.05 (5.72)
79.25 (5.66)
77.78 (5.36)
89.08 (4.26)
100 (0.00)
67.54 (7.08)
69.66 (6.75)
69.40 (6.37)
82.95 (5.70)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.12 (0.325)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.48 (0.14)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.21 (0.22)
2.00 (0.00)
1.48 (0.12)
1.49 (0.13)
12
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
70.39 (3.82)
74.50 (4.00)
73.35 (3.49)
85.06 (5.19)
100 (0.00)
62.78 (5.20)
64.03 (4.99)
63.53 (4.78)
76.71 (6.47)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.14 (0.35)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.46 (0.14)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.26 (0.13)
2.00 (0.00)
1.50 (0.11)
1.52 (0.10)
Figure 10: Effects of dimensionality. Training size = 100. No noise. Ground truth tress are size 2.
The third set of experiments examines the effect of problem dimensionality while keeping the tree size ($K=2$) and training set size ($N=100$) constant. ConTree and CART-depth were fixed at depth 2, and CART-size was fixed at size 4. Table 10 shows the effect of increasing the number of features for a fixed training size and tree size.
Although HODT might be expected to suffer most from the increased combinatorial complexity associated with higher dimensionality, it remains the most robust method on both training and test datasets, exhibiting the smallest decrease in accuracy. Increasing the number of features significantly affects all methods: CART-depth experiences decreases of approximately 18.61% in training accuracy and 21.84% in test accuracy; CART-size decreases by 19.99% (train) and 25.54% (test); ConTree decreases by 19.52% (train) and 24.98% (test); whereas HODT decreases only by 14.88% (train) to 21.26% (test). This robustness may be attributed to the effectiveness of $\mathit{hodtCoreset}$, which efficiently explores configurations without being heavily affected by combinatorial complexity.
Interestingly, our results for axis-parallel ODT algorithms on datasets with hyperplane decision tree ground truth differ from those reported by Bertsimas and Dunn (2017). However, it is important to note that our experimental setup differs from theirs. Bertsimas and Dunn (2017) observed that the performance gap between axis-parallel ODT algorithms and approximate CART increases in higher dimensions, with little difference in lower dimensions. In contrast, our experiments show that at lower dimensions, there is a significant difference between CART-depth and ConTree (3.89%), which decreases to 0.75% when $D=12$.
Effect of noise on labels
Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
Maximum Average Expected
0
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.42 (4.61)
95.19 (3.11)
93.49 (3.60)
99.62 (0.68)
1 (0.00)
87.25 (5.20)
91.21 (3.94)
89.99 (3.71)
98.06 (1.21)
1 (0.00)
2.92 (0.27)
4.0 (0.00)
2.98 (0.04)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.97 (0.09)
2.49 (0.16)
1.99 (0.05)
1.67 (0.00)
1.67 (0.00)
1.99 (0.03)
2.11 (0.23)
1.98 (0.11)
1.67 (0.28)
1.68 (0.28)
5
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
86.29 (4.48)
90.64 (3.07)
89.07 (3.45)
95.00 (0.20)
0.95 (0.00)
87.15 (4.89)
91.01 (3.67)
89.87 (3.47)
97.20 (1.40)
1 (0.00)
2.94 (0.24)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.2.00 (0.40)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.98 (0.08)
2.50 (0.16)
2.00 (0.00)
1.68 (0.00)
1.67 (0.00)
1.99 (0.02)
2.10 (0.24)
2.00 (0.00)
1.62 (0.29)
1.68 (0.28)
10
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
82.25 (3.78)
86.25 (2.67)
84.69 (3.09)
90.45 (0.49)
0.90 (0.00)
86.74 (5.36)
90.45 (4.07)
89.48 (3.96)
97.59 (1.21)
100 (0.00)
2.98 (0.14)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.12 (0.32)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
3.12 (0.32)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.16 (0.26)
2.00 (0.00)
1.64 (0.30)
1.68 (0.28)
15
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
77.90 (3.99)
81.80 (3.10)
80.53 (3.04)
85.69 (0.54)
84.99 (0.00)
85.99 (5.71)
90.19 (3.67)
89.10 (3.79)
97.10 (1.84)
100 (0.00)
2.98 (0.14)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.22 (0.41)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.51 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.04)
2.19 (0.31)
2.00 (0.00)
1.63 (0.29)
1.68 (0.28)
20
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
73.88 (3.16)
77.36 (2.91)
76.28 (2.81)
80.98 (0.69)
79.90 (0.00)
86.06 (4.95)
89.51 (4.59)
88.61 (4.43)
96.58 (2.06)
100 (0.00)
2.94 (0.24)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.98 (0.08)
2.54 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.13 (0.34)
2.00 (0.00)
1.68 (0.27)
1.68 (0.28)
25
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
70.35 (3.22)
73.19 (2.54)
72.26 (2.62)
76.37 (0.83)
75.00 (0.00)
84.59 (6.66)
87.78 (5.41)
87.06 (5.16)
95.67 (2.93)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.26 (0.44)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.52 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.24 (0.32)
2.00 (0.00)
1.67 (0.78)
1.68 (0.28)
Table 3: The effect of noise on labels. Training data size = 100. Ground truth trees are size 2.
In the fourth set of experiments, we introduced noise in the labels. As the noise level increased, it was noteworthy that HODT was able to find solutions that outperformed the ground truth on the training dataset. As before, we fixed $D=2$ and ($K=2$). Following the experimental setup of Bertsimas and Dunn (2017) we added noise by increasing the label of a random $k\%$ of the points by 1, where Table 3 presents the results.
As the noise level increased, the accuracy of all methods tended to decrease, with similar effects on out-of-sample performance up to a noise level of 20%. Beyond this point (from 20% to 25% noise), differences became more pronounced: CART-depth, CART-size, and ConTree decreased by 1.47%, 1.73%, and 1.55%, respectively, whereas HODT decreased by only 0.91%. From a broader perspective, when increasing noise from 0% to 25%, HODT’s out-of-sample accuracy decreased by 2.39%, compared with 2.93%, 3.43%, and 2.66% for ConTree, CART-size, and CART-depth, respectively. Notably, starting at a 10% noise level, HODT was able to find solutions that exceeded the ground truth on the training dataset while maintaining strong performance on out-of-sample tests. These results further refute the notion that optimal methods are less robust to noise, highlighting the superiority of HODT.
Effect of noise on data
Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
Maximum Average Expected
0
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.29 (4.74)
94.20 (3.70)
92.73 (3.90)
99.88 (0.26)
100 (0.00)
87.23 (5.14)
90.32 (4.24)
89.55 (4.10)
98.08 (1.00)
100 (0.00)
2.96 (0.20)
4.0 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.2 (0.399)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.065)
2.51 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.09 (0.33)
2.00 (0.00)
1.63 (0.29)
1.63 (0.28)
5
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.16 (4.63)
93.98 (3.78)
92.62 (3.79)
99.62 (0.40)
99.59 (0.44)
87.16 (5.06)
90.32 (4.20)
89.50 (3.94)
98.10 (1.01)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.05)
2.53 (0.168)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.01)
2.05 (0.32)
2.00 (0.00)
1.64 (0.28)
1.65 (0.28)
10
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
90.30 (4.55)
93.98 (3.57)
92.50 (3.65)
99.07 (0.62)
98.81 (0.71)
87.36 (5.09)
90.30 (4.55)
89.38 (4.09)
97.50 (0.90)
100 (0.00)
2.96 (0.20)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.28 (0.45)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.065)
2.54 (0.17)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.03 (0.34)
2.00 (0.00)
1.58 (0.29)
1.65 (0.28)
15
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.86 (4.91)
93.78 (3.92)
92.52 (3.83)
99.05 (0.66)
98.72 (0.80)
87.06 (4.97)
90.27 (4.42)
89.31 (4.38)
97.90 (1.10)
100 (0.00)
2.96 (0.20)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.32 (0.47)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.07)
2.55 (0.18)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.02)
2.03 (0.33)
2.00 (0.00)
1.63 (0.28)
1.65 (0.28)
20
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.75 (4.98)
93.75 (3.70)
92.39 (3.72)
98.65 (0.87)
98.11 (1.04)
86.61 (5.57)
90.21 (4.27)
89.25 (4.05)
97.40 (11.50)
100 (0.00)
3.00 (0.00)
4.00 (0.00)
3.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.18 (0.38)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
2.49 (0.16)
2.00 (0.00)
1.67 (0.00)
1.67 (0.00)
2.00 (0.00)
2.08 (0.27)
2.00 (0.00)
1.64 (0.28)
1.65 (0.28)
25
CART-depth-2
CART-size-4
ConTree-2
HODT
Ground Truth
89.76 (4.47)
93.58 (4.13)
92.29 (3.65)
98.35 (1.00)
97.67 (0.91)
87.13 (4.93)
90.28 (4.50)
89.55 (4.17)
97.62 (1.20)
100 (0.00)
2.98 (0.14)
4.00 (0.00)
2.98 (0.00)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
3.24 (0.43)
2.00 (0.00)
2.00 (0.00)
2.00 (0.00)
1.99 (0.046)
2.53 (0.17)
1.99 (0.05)
1.67 (0.00)
1.67 (0.00)
1.99 (0.09)
2.05 (0.35)
1.99 (0.05)
1.67 (0.27)
1.65 (0.28)
Figure 11: The effect of noise on training data. Training data size = 100. Ground truth trees are size 2.
Finally, in the last set of experiments, we examined the effect of noise in the features of the data. We fixed $N=100$, $D=2$ and $K=2$. Again, HODT was able to find solutions that outperformed the ground truth on the training dataset while achieving the best out-of-sample performance. Combined with previous experiments, these results strongly refute the idea that optimal algorithms necessarily overfit the data. Even when training accuracy exceeds that of the ground truth, proper control of model complexity prevents overfitting.
In summary, across all synthetic data experiments, HODT-generated trees most closely matched the quality metrics of the ground truth trees. HODT not only produced more accurate results but also generated smaller trees, demonstrating a clear advantage in scenarios where tree size must be strictly controlled. Interestingly, by allowing slightly more flexibility in size or depth, CART can sometimes achieve slightly better solutions than optimal axis-parallel decision tree algorithms. However, in contexts requiring strict control of tree size, HODT consistently provides superior performance compared to all other methods.
Moreover, our experiments align with the observations of Bertsimas and Dunn (2017), countering the widely held misconception that optimal methods are more prone to overfitting the training set at the expense of out-of-sample accuracy. We demonstrate that optimal solutions remain robust in data-scarce and noisy settings, even when training accuracy exceeds that of the ground truth.
5.4 Computational experiments on real-world datasets
Dataset NN $D$ $C$ CART-depth $d=2$ ConTree $d=2$ CART-size $K=2$ CART-size $K=3$ HODT $K=2$ HODT $K=3$
haberman 283 3 2 75.58/73.68 (1.62/3.14) 76.991/72.98 (0.93/3.06) 74.25/72.63 (1.26/4.52) 75.31/73.68 (1.82/2.94) 79.38/78.25 (1.23/3.60) 80.70/78.24 (1.39/4.87)
BldTrns 502 4 2 75.71/72.08 (1.05/2.02 77.16/71.09 (0.60/01.70) 75.711/72.08 (1.05/02.02) 76.91/74.06 (0.56/03.09) 79.38/78.50 (1.16/4.37) 80.71/78.25 (1.12/3.86)
spesis 975 3 2 94.13/93.64 (0.30/1.06) 94.26/93.74 (0.30/0.88) 94.13/93.64 (0.30/1.06) 94.18/93.64 (0.35/1.06) 94.97/93.85 (0.44/1.95) 95.39/93.85 (0.43/1.95)
algerian 243 14 2 99.28/98.37 (0.25/1.53) 99.69/95.92 (0.25/2.24) 99.28/98.37 (0.25/1.52) 99.59/98.37 (0.21/1.53) 99.59/97.96 (0.43/1.44) 100/98.78 (0.00/1.12)
Cryotherapy 89 6 2 94.93/90.00 (1.91/6.48) 94.93/91.11 (1.91/6.67) 91.83/80.00 (2.87/10.30) 94.93/90.00 (1.91/6.48) 98.31/93.33 (0.63/6.33) 99.16/93.33 (0.77/4.65)
Caesarian 72 5 2 73.33/58.67 (2.05/7.78) 75.09/57.33 (1.31/6.80) 73.33/6133 (2.05/7.78) 75.79/64.00 (2.05/09.04) 88.42/85.33 (1.57/5.58) 91.23/84.00 (2.15/3.68)
ecoli 336 7 8 80.37/79.12 (1.28/2.35) 81.11/80.59 (0.80/3.65) 76.04/75.00 (1.38/2.94) 81.11/77.35 (0.80/1.77) 80.73/77.64 (0.97/4.08) 82.76/79.12 (1.10/3.81)
GlsId 213 9 6 62.35/62.33 (1.18/5.38) 67.29/62.79 (1.15/04.88) 61.76/62.33 (1.18/5.38) 66.47/63.72 (1.12/5.22) 71.18/61.86 (2.46/4.53) 74.25/62.79 (2.29/5.20)
Parkinsons 195 22 2 87.95/84.62 (1.48/5.85) 93.20/82.56 (1.04/5.23) 87.82/85.13 (1.67/6.36 90.77/86.67 (1.04/5.71 93.72/89.74 (1.66/3.14) 94.49/89.74 (2.15/3.14
Diabetic 1146 19 2 65.55/64.00 (0.92/3.74) 67.36/63.48 (0.79/3.95) 65.55/64.00 (0.92/3.74) 67.53/62.96 (1.05/2.66) 79.93/76.52 (1.17/1.71) 80.09/76.35 (1.06/2.38)
BalScl 625 4 3 71.72/65.12 (0.94/2.00) 72.88/66.88 (0.35/1.40) 69.80/63.20 (0.44/1.75) 72.68/64.64 (1.12/2.74) 93.44/94.40 (0.33/1.26) 94.28/94.56 (0.30/1.54)
StatlogVS 845 18 4 53.34/52.66 (0.32/2.97) 62.42/61.89 (0.91/4.07) 49.94/48.40 (0.64/3.73) 53.34/52.66 (0.32/2.97) 65.12/63.31 (0.78/4.73) 68.75/64.38 (0.74/2.58)
ImgSeg 210 19 7 50.00/41.91 (6.42/9.71) 58.69/49.04 (0.80/3.23) 44.64/33.81 (1.30/5.51) 57.86/52.38 (1.09/4.26) 66.00/41.34 (2.58/6.24) 66.47/49.52 (1.91/5.20)
iris 147 4 3 97.09/90.67 (0.68/2.49) 97.09/90.66 (0.68/2.49) 97.09/90.67 (0.68/2.49) 98.12/91.33 (1.26/3.40) 99.80/97.33 (3.49/7.14) 98.97/96.77 (4.18/6.31)
MnkPrb 432 6 2 74.44/77.24 (1.53/6.06) 78.38/75.40 (1.49/5.89) 74.44/77.24 (1.53/6.06) 81.39/82.76 (3.28/6.86) 82.24/79.75 (1.56/3.96) 82.26/86.20 (2.79/2.44)
UKM 403 5 5 81.18/75.80 (1.52/4.18) 82.17/76.05 (0.54/1.84) 78.14/75.31 (0.72/4.18) 82.05/75.80 (0.50/2.59) 71.24/64.69 (9.41/16.69) 72.42/68.15 (10.74/16.84)
TchAsst 106 5 4 54.76/39.09 (1.68/5.46) 57.14/41.81 (1.30/5.30) 53.33/40.00 (0.48/4.45) 56.43/42.73 (2.33/4.64) 71.19/57.27 (10.0/7.61) 75.71/60.90 (1.60/9.96)
RiceCammeo 3810 7 2 92.91/92.52 (0.15/0.67) 93.30/92.91 (0.19/0.80) 92.91/92.52 (0.15/0.67) 92.91/92.52 (0.15/0.67) 94.00/93.28 (0.29/1.28) 94.18/93.31 (0.29/1.39)
Yeast 1453 8 10 48.80/45.77 (0.35/0.83) 49.81/49.28 (0.18/01.55) 46.56/44.19 (0.43/0.77) 48.80/45.77 (0.35/0.83) 48.62/45.66 (0.66/1.38) 49.60/45.57 (0.70/3.09)
WineQuality 5318 11 7 52.96/52.88 (0.39/0.45) 54.07/53.85 (0.27/1.07) 52.45/52.59 (0.93/0.58) 52.96/52.88 (0.39/0.45) 54.70/55.23 (0.31/0.98) 55.09/55.23 (0.29/1.16)
SteelOthers 1941 27 2 70.59/71.11 (0.21/1.32) 73.75/71.31 (0.31/0.85) 70.50/71.05 (0.36/1.26) 72.35/72.55 (0.89/1.21) 77.43/76.52 (0.05/0.81) 78.09/75.83 (1.15/1.98)
DrgCnsAlc 1885 12 7 69.07/69.87 (0.48/1.92) 69.54/69.66 (0.39/1.90) 69.07/69.87 (0.48/1.92) 69.07/69.87 (0.48/1.92) 70.89/69.07 (0.35/1.33) 71.70/69.23 (0.36/1.48)
DrgCnsImp 1885 12 7 40.41/39.79 (0.57/2.03) 41.95/38.04 (0.39/0.87) 40.36/39.89 (0.55/2.19) 40.41/39.79 (0.57/2.03) 45.04/42.55 (0.88/3.07) 45.73/42.60 (2.39/1.80)
DrgCnsSS 1885 12 7 51.54/51.88 (0.36/2.75) 52.73/52.20 (0.35/1.70) 51.53/52.79 (0.35/1.38) 51.54/51.88 (0.36/2.75) 54.13/52.31 (0.58/1.70) 54.93/52.10 (0.51/1.28)
EstObLvl 2087 16 7 55.36/55.98 (0.18/0.66) 55.40/55.31 (0.14/1.13) 43.14/42.39 (0.30/1.33) 55.36/55.98 (0.18/0.66) 46.05/44.21 (1.09/2.87) 51.99/50.28 (1.82/2.64)
AiDS 2139 23 2 85.44/85.70 (0.39/01.43) 87.54/86.73 (0.11/0.58) 85.44/85.70 (0.39/1.43) 89.18/88.51 (0.20/0.68) 86.90/85.79 (1.24/1.93) 87.13/85.65 (1.31/1.94)
AucVer 2043 7 2 90.42/89.58 (0.55/1.57) 91.00/89.98 (0.40/1.59) 90.27/89.49 (0.40/1.61) 90.42/89.58 (0.55/1.57) 89.65/88.46 (0.35/1.99) 89.87/88.66 (0.34/1.83)
Ai4iMF 10000 6 2 97.25/97.07 (0.11/0.36) 97.39/97.31 0.07/0.26) 97.08/97.00 00.27/00.34) 97.16/97.03 0.22/0.36) 97.79/97.95 0.11/0.18) 97.90/98.00 0.12/0.15)
VoicePath 704 2 2 96.87/95.60 (0.51/2.12) 97.16/96.31 (0.56/2.43) 96.87/95.60 (0.51/2.12) 97.16/95.04 (0.48/1.85) 97.69/97.44 (0.28/1.08) 97.90/97.31 (0.32/1.17)
WaveForm 5000 21 3 70.69/68.50 (0.25/1.33) 71.28/68.54 (0.34/0.81) 66.31/65.02 (0.74/0.87) 70.69/68.50 (0.25/1.33) 67.05/66.62 (2.44/1.55) 65.12/65.80 (1.82/1.65)
Table 4: Five-fold cross-validation results on the UCI dataset. We compare the performance of our HODT algorithm, with $K$ (number of splitting rules) ranging from 2 to 3, trained using $\mathit{sodtWSH}$—against approximate methods:size- and depth-constrained CART algorithms (CART-size and CART-depth), as well as the state-of-the-art optimal axis-parallel decision tree algorithm, ConTree. The depth of the CART-depth and ConTree algorithms are fixed at 2. Results are reported as mean 0-1 loss on the training and test sets in the format Training Error / Test Error (Standard Deviation: Train / Test). The best-performing algorithm in each row is shown in bold.
We now present a direct comparison between HODT, ConTree, CART-depth, and CART-size on real-world datasets. Our goal is to evaluate the effectiveness of the more flexible hyperplane decision tree model relative to the standard axis-parallel decision tree models. Table 4 reports the mean out-of-sample accuracy for 30 classification datasets from the UCI Machine Learning Repository.
As expected, when model complexity is controlled, HODT outperforms the axis-parallel methods on most datasets. Specifically, HODT achieves higher training accuracy on 25 datasets and better out-of-sample accuracy on 23 datasets. By contrast, ConTree achieves the best training accuracy on only 4 datasets and the best out-of-sample accuracy on 6 datasets. The CART-depth algorithm achieves the highest out-of-sample accuracy on just 3 datasets, performing clearly worse than the optimal ConTree under the same constraints. Even with greater flexibility (allowing tree depths beyond 2), CART-depth achieves the highest accuracy on only 5 datasets, still falling behind HODT.
In summary, the more flexible hyperplane decision tree model shows strong potential for interpretable learning tasks under controlled model complexity, often delivering substantially better performance than axis-parallel models. For instance, on the BalScl dataset, HODT improves training accuracy by over 20% and test accuracy by nearly 30% compared to the optimal axis-parallel tree algorithm, ConTree. These results underscore the benefits of moving beyond axis-parallel and even hyperplane-based approaches, demonstrating that hypersurface decision trees can capture richer and more complex decision boundaries. Overall, our findings indicate that HODT not only achieves higher accuracy but also offers a more general and robust framework for decision tree construction, paving the way for broader applications in machine learning where interpretability and predictive power are equally critical.
6 Conclusion and future directionsa
Conclusion
In this second of two papers, we first identified three types of ancestry relations between hypersurfaces—mutual ancestry, asymmetrical ancestry, and no ancestry (i.e., crossed hyperplanes). We then developed an incremental method to efficiently construct the ancestry relation matrix by leveraging the prefix-closed property of crossed hyperplanes, thereby eliminating any combinations of ancestry relations that contain a pair of crossed hyperplanes. Building on this foundation, we proposed the first algorithm for solving the optimal decision tree problem with hypersurface splits based on the algorithmic and geometric principles established in the first part. Due to the intractable combinatorics of the hypersurface decision problem, our algorithm cannot obtain exact solutions for most datasets, even for modest sizes. To evaluate the empirical performance of hyperplane decision tree models, we proposed two heuristic methods for training them.
In our experiments on synthetic datasets, we generated data such that the ground truth was given by random hyperplane decision trees. Our results showed that hyperplane decision tree models are substantially more accurate than axis-parallel decision tree algorithms when the ground truth consists of convex polygons generated by hyperplane splits. Moreover, even when a model achieves higher training accuracy than the ground truth in noisy settings, its out-of-sample accuracy remains extremely high, refuting the idea that optimal models necessarily overfit the data. Experiments on real-world datasets further demonstrate the superiority of hyperplane decision tree models when model complexity is properly controlled.
Future research directions
There remain several limitations in our current research that warrant further investigation.
First, implementation of additional algorithms. In our experiments, we implemented only $\mathit{odt}_{\text{rec}}$ and $\mathit{odt}_{\text{vec}}$, while $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ were not considered. We note that $\mathit{odt}_{\text{depth}}$ is challenging to parallelize compared with $\mathit{odt}_{\text{size}}$ , $\mathit{odt}_{\text{kperms}}$ and shares features with $\mathit{odt}_{\text{vec}}$. Nevertheless, both $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ has their advantages: $\mathit{odt}_{\text{depth}}$ has advantages for constructing large trees efficiently using a depth-first search strategy. It would be interesting to explore whether $\mathit{odt}_{\text{depth}}$ could be as effective for hyperplane decision trees as it is for axis-parallel decision trees, as reported in Brița et al. (2025).
Second, improved heuristics. The current two heuristics become ineffective when $K$ and $D$ are large, providing only marginal improvements or even deteriorating accuracy. Randomly generating hyperplanes is ineffective because the probability of finding feasible nested combinations is low. Future work could focus on developing more powerful heuristics for constructing hyperplane decision trees in high-dimensional or large-$K$ settings. One possible approach is to combine depth-first search with coreset selection methods.
Third, conflicts between the general position assumption and categorical data. The correctness of $\mathit{hodt}$ relies on the general position assumption for data points. However, this assumption is reasonable primarily for numerical classification tasks, and real-world datasets often violate it—for example, categorical datasets with discrete features. In such cases, axis-parallel decision tree algorithms may be more appropriate than hyperplane decision tree algorithm. For categorical data, rather than creating $D$ splits for each data point, the algorithms of Brița et al. (2025) and Mazumder et al. (2022) efficiently exploit the structure of categorical features by eliminating duplicate splitting rules from the candidate sets. However, it remains unclear how to leverage categorical information in the hyperplane or hypersurface setting. Investigating ways to incorporate this assumption into hypersurface decision tree algorithms could lead to more efficient methods for categorical datasets.
Fourth, algorithms with mixed splitting rules. In Part I of this paper, we proposed a generator for mixed-splitting rules. In principle, this generator can be used as input for $\mathit{odt}$ to produce decision tree with mixed-splitting rules; however, due to space constraints, we did not perform experiments with mixed-splitting rules. Future work could explore the construction of potentially more flexible models using these rules.
Finally, extensions to random forests. Two key results from Breiman (2001) suggest promising directions for constructing random forests with hypersurface decision trees.shows that adding more trees does not lead to overfitting but converges to a limiting generalization error, while Theorem 2.3 establishes that generalization performance improves with the strength of individual classifiers. These insights align naturally with our framework: even in our Python implementation, the $\mathit{sodt}$ program can efficiently solve $1\times 10^{6}$ feasible nested combinations in tens of seconds. Moreover, flexible tree models often achieve stronger performance than axis-parallel trees. Future research could explore constructing large random forests with hypersurface decision trees, potentially yielding ensembles that outperform those based on axis-parallel trees and offering a promising direction for further study.
Appendix A Algorithms
A.1 nested-combination generator
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
  2. 2.
    Output: Array of $\left(K,G\right)$-nested-combinations
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  8. 8.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  9. 9.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  10. 10.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  11. 11.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  12. 12.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  13. 13.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  14. 14.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  15. 15.
    return $\mathit{ncss}\left[K\right]$
Algorithm 5 $\mathit{nestedCombs}$
A.2 Incremental ancestry relation matrix generator
  1. 1.
    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
  2. 2.
    Output: Array of $\left(K,G\right)$-nested-combinations
  3. 3.
    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
  4. 4.
    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
  5. 5.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
  6. 6.
    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
  7. 7.
    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
  8. 8.
      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
  9. 9.
      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
  10. 10.
    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
  11. 11.
    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
  12. 12.
    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
  13. 13.
       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
  14. 14.
        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
  15. 15.
        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // $\mathit{update}_{\text{arMat}}$ update $\left(k,G\right)$-nested-combinations and its associated ancestry relation matrix, defined in Algorithm 1 for hyperplane spliting rules.
  16. 16.
    return $\mathit{ncss}\left[K\right]$
Algorithm 6 $\mathit{\mathit{\mathit{nestedCombsFA}}}\left(K,G,\mathit{xs}\right)$
A.3 Experiments for $K>3$ decision tree
Dataset NN $D$ $C$ CART-depth $d=3$ ConTree $d=3$ CART-size $K=4$ CART-size $K=5$ CART-size $K=6$ HODT $K=4$ HODT $K=5$ HODT $K=6$
haberman 283 3 2 77.35/74.39 (1.38/3.25) 80.97/72.98 (1.22/03.94) 76.20/73.33 (1.99/2.58) 77.17/72.28 (2.21/1.72) 77.97/72.28 (1.71/2.58) 81.68/78.94 (0.86/3.51) 82.48/79.29 (0.97/4.30) 82.83/79.65 (1.15/4.37)
BldTrns 502 4 2 77.41/74.06 (0.64/03.62) 80.45/74.06 (0.843/2.76) 78.40/77.23 (0.56/2.80) 79.05/75.45 (1.32/4.13) 79.20/73.86 (1.21/3.41) 81.68/78.94 (0.97/3.51) 82.48/79.30 (1.20/2.88) 82.83/79.65 (1.15/4.23)
spesis 975 3 2 94.23/93.74 (0.35/1.19) 94.56/93.44 (0.34/1.10) 94.28/93.54 (0.33/1.11) 94.39/93.54 (0.34/0.95) 94.44/93.54 (0.29/0.95) 95.77/93.85 (0.43/1.95) 96.10/93.85 (0.43/1.95) 96.44/94.05 (0.46/1.72)
algerian 243 14 2 99.59/98.37 (0.21/1.53) 100/95.51 (0.00/3.00) 99.79/98.37 (0.25/1.53) 99.90/98.37 (0.21/1.53) 100/98.37 (0.00/1.53) 100/99.18 (0.00/1.12 100/97.96 (0.00/1.44) 100/96.74 (0.00/1.83)
Cryotherapy 89 6 2 94.93/90.00 (1.91/06.48) 99.44/78.89 (0.69/12.86) 94.93/90.00 (1.91/6.48) 96.62/86.67 (1.91/9.03) 97.75/86.67 (1.69/9.03) 99.44/94.44 (0.77/3.93) 99.72/93.33 (0.77/3.93) 99.72/92.22 (0.49/1.95)
Caesarian 72 5 2 77.90/58.67 (1.789/7.78) 82.11/58.67 (1.72/4.99) 77.54/66.67 (2.33/4.22) 78.60/56.00 (3.58/12.36) 80.35/57.33 (2.58/6.80) 92.63/85.33 (1.47/2.98) 93.68/86.68 (20.96/0.00) 93.68/85.33 (20.96/2.98)
ecoli 336 7 8 85.52/81.47 (0.37/1.77) 87.84/82.06 (0.18/2.85) 85.07/81.76 (0.53/1.50) 85.82/82.65 (0.88/2.53) 85.97/82.94 (0.65/2.20) 84.25/79.12 (1.70/4.08) 84.40/79.41 (21.25/4.29 85.90/79.12 (21.25/4.46)
GlsId 213 9 6 72.35/66.05 (1.86/6.51) 80.12/70.70 (0.941/3.78) 71.29/66.98 (1.72/6.14) 74.24/63.72 (1.46/4.79) 75.88/64.65 (2.10/7.98) 76.35/61.86 (1.97/4.53) 77.77/61.86 (21.92/4.53) 77.29/58.61 (21.64/4.47)
Parkinsons 195 22 2 95.00/87.18 (10.26/66.86) 99.35/89.23 (0.57/4.97) 91.79/85.64 (1.74/6.20) 94.74/86.67 (0.85/7.50) 96.15/87.69 (0.70/8.17) 94.49/87.69 (2.15/4.22) 95.87/89.23 (21.23/3.89) 97.82/88.76 (21.36/2.55)
Diabetic 1146 19 2 67.97/63.65 (1.27/2.31) 71.63/65.39 (0.81/3.43) 67.99/64.43 (1.02/2.54) 68.71/63.91 (0.74/2.76) 69.30/64.17 (0.57/1.66) 80.13/76.35 (1.44/2.89) 80.09/76.43 (21.30/2.22) 80.78/74.87 (20.86/1.72)
BalScl 625 4 3 76.96/70.08 (0.71/3.60) 78.48/72.3 (0.64/3.97) 75.28/67.68 (0.37/4.00) 76.92/69.76 (1.15/4.12) 77.12/70.88 (1.31/5.00) 95.08/94.49 (0.44/1.34) 95.60/94.72 (20.40/1.84) 95.84/94.88 (20.36/2.01)
StatlogVS 845 18 4 67.84/64.73 (2.20/5.29) 72.01/64.85 (0.96/2.66) 63.76/62.49 (1.02/5.18) 65.33/62.72 (0.68/5.55) 66.07/63.91 (0.78/4.73) 66.83/63.31 (0.90/2.35) 66.39/63.07 (20.93/1.28) 65.21/61.23 (21.86/3.76)
ImgSeg 210 19 7 63.33/58.10 (6.70/6.67) 88.21/80.95 (1.02/5.83) 70.95/64.76 (1.15/3.81) 82.38/77.62 (1.44/5.75) 88.69/82.38 (1.30/4.90) 69.17/48.57 (1.60/5.58) 69.17/48.57 (24.04/5.16) 69.40/47.14 (1.48/6.52)
iris 147 4 3 98.12/90.67 (1.26/2.49) 99.82/91.33 (0.34/4.00) 98.80/88.00 (1.03/4.52) 99.15/88.00 (0.76/4.52) 99.32/88.00 (0.64/4.52) 99.15/96.77 (2.94/12.61) 100/97.77 (0.00/12.61) 100/96.77 (0.00/12.61)
MnkPrb 432 6 2 82.44/78.62 (3.61/3.89) 89.51/86.44 (1.43/5.65) 82.43/78.62 (3.61/3.89) 83.25/78.16 (2.41/2.72) 81.54/77.01 (2.60/3.17) 82.46/81.38 (1.30/5.6) 82.73/81.61 (1.61/5.75) 82.96/82.30 (1.50/6.38)
UKM 403 5 5 88.76/86.42 (1.38/3.75) 91.55/87.90 (0.60/3.26) 85.34/80.00 (0.46/3.86) 87.39/83.70 (0.64/3.44) 89.57/87.65 (0.32/1.35) 73.10/68.89 (10.74/17.00) 73.60/75.30 (10.17/9.88) 73.79/75.56 (10.3/9.54)
TchAsst 106 5 4 62.381/41.82 (2.21/4.45) 67.62/46.36 (2.65/10.52) 58.81/40.91 (1.43/4.98) 62.14/43.64 (2.31/3.64) 63.81/42.73 (2.88/4.64) 78.33/62.73 (1.13/10.85) 79.76/62.73 (1.19/10.85) 80.48/59.09 (1.81/8.50)
RiceCammeo 3810 7 2 93.06/92.49 (0.28/0.79) 93.79/92.36 (0.16/0.89) 92.91/92.52 (0.15/0.67) 93.03/92.49 (0.28/0.72) 93.09/92.49 (0.31/0.76) 94.27/93.49 (0.29/1.36) 94.35/93.57 (0.29/1.31) 94.38/93.57 (0.31/1.20)
Yeast 1453 8 10 57.86/54.35 (0.62/0.98) 58.50/55.74 (0.38/0.96) 55.83/52.16 (0.31/1.78) 56.99/53.13 (0.42/0.89) 57.95/54.36 (0.68/1.20) 50.02/46.83 (0.92/2.73) 51.23/47.23 (1.02/3.01) 50.23/48.11 (1.03/2.50)
WineQuality 5318 11 7 53.95/53.29 (0.40/0.57) 55.52/53.29 (0.19/0.81) 52.96/52.88 (0.39/0.45) 53.09/52.84 (0.33/0.46) 53.35/53.08 (0.41/0.61) 55.21/55.60 (0.37/1.24) 54.96/55.02 (0.32/0.98) 55.42/54.32 (1.21/0.42)
SteelOthers 1941 27 2 73.30/73.57 (1.09/0.44) 77.77/73.57 (0.16/1.33) 73.17/73.37 (0.82/1.06) 74.38/74.50 (1.55/1.02) 76.10/75.32 (0.96/0.88) 76.49/75.42 (0.81/0.74) 77.81/74.38 (0.32/0.56) 74.32/72.5 (0.34/1.83)
DrgCnsAlc 1885 12 7 69.35/69.44 (0.43/1.91) 70.623/69.55 (0.50/1.85) 69.24/69.71 (0.69/2.12) 69.44/69.34 (0.55/2.02) 70.03/70.08 (0.78/1.74) 72.44/69.18 (0.34/1.45) 72.88/68.96 (0.27/1.42) 73.21/69.43 (0.43/1.58)
DrgCnsImp 1885 12 7 41.22/38.57 (0.52/1.38) 43.83/38.20 (0.43/3.13) 40.69/39.47 (0.67/1.87) 41.07/38.99 (0.69/1.61) 41.27/38.89 (0.73/1.59) 46.18/42.65 (2.43/4.38) 46.25/42.60 (3.18/4.37) 46.06/40.80 (4.01/2.32)
DrgCnsSS 1885 12 7 52.06/51.03 (0.53/0.024) 54.23/51.57 (0.38/2.22) 51.54/51.88 (0.36/2.75) 51.80/51.41 (0.22/2.39) 51.80/51.41 (0.22/2.39) 55.48/51.57 (0.50/1.60) 55.82/51.36 (0.60/1.33) 56.21/52.42) (0.32/1.24
EstObLvl 2087 16 7 65.07/63.16 (0.12/2.07) 72.51/70.05 (0.28/1.54) 61.25/60.62 (0.35/1.40) 67.31/67.22 (0.31/1.33) 69.50/68.66 (0.23/1.4) 51.70/48.95 (1.61/3.17) 49.61/46.89 (1.75/2.51) 49.64/49.22 (1.46/2.84)
AiDS 2139 23 2 89.23/88.55 (0.15/0.74) 90.24/89.11 (0.20/0.86) 89.23/88.55 (0.15/0.74) 89.36/88.60 (0.27/0.80) 89.49/88.64 (0.23/0.84) 86.99/85.93 (1.87/2.23) 86.99/85.65 (2.19/2.51) 87.02/85.23 (1.32/2.44)
AucVer 2043 7 2 92.08/90.56 (0.65/2.22) 94.98/93.55 (0.40/1.50) 91.91/90.27 (0.64/2.16) 92.83/91.00 (0.52/2.08) 93.05/91.30 (0.46/2.21) 90.00/88.36 (0.28/1.53) 90.11/88.66 (0.48/1.38) 89.92/88.31 (0.35/1.65)
Ai4iMF 10000 6 2 97.41/97.23 (0.05/0.48) 97.84/97.27 (0.06/0.22) 97.22/97.06 (0.28/0.39) 97.49/97.19 (0.26/0.32) 97.56/97.26 (0.19/0.38) 97.92/98.01 (0.11/0.17) 97.88/97.92 (0.12/0.17) 97.82/97.85 (0.17/0.09)
VoicePath 704 2 2 97.16/95.04 (0.48/1.85) 98.15/96.03 (0.43/1.83) 97.41/95.04 (0.53/2.24) 97.58/95.18 (0.47/2.26) 97.73/95.04 (0.51/2.33) 98.12/97.89 (0.27/1.42) 98.30/97.89 (0.27/1.00) 98.40/98.01 (0.25/1.27)
WaveForm 5000 21 3 73.29/70.44 (0.32/0.41) 76.53/73.26 (0.16/1.03) 71.37/68.84 (0.45/1.17) 72.36/69.84 (0.62/1.16) 72.69/69.96 (0.79/1.17) 64.37/62.70 (0.82/1.11) 64.58/62.70 (3.71/1.82) 66.12/62.70 (2.52/3.64)
Table 5: Five-fold cross-validation results on the UCI dataset. We compare the performance of our HODT algorithm, with $K$ (number of splitting rules) ranging from 4 to 6, trained using $\mathit{sodtWSH}$—against approximate methods:size- and depth-constrained CART algorithms (CART-size and CART-depth), as well as the state-of-the-art optimal axis-parallel decision tree algorithm, ConTree. The depth of the CART-depth and ConTree algorithms are fixed at 3. Results are reported as mean 0-1 loss on the training and test sets in the format Training Error / Test Error (Standard Deviation: Train / Test). The best-performing algorithm in each row is shown in bold.

### v2

@@
 We also analyzed generalization performance across 30 real-world datasets, showing that HODT can achieve up to 30% higher accuracy than the state-of-the-art optimal axis-parallel decision tree algorithm when tree complexity is properly controlled.
-(a) Ground truth, 0-1 loss: 0, KK=5, dd=3
-(b) CART, 0-1 loss: 56, KK=15, dd=5
-(c) ConTree, 0-1 loss: 34, KK=15, dd=4
-(d) HODT (Our), 0-1 loss: 18, KK=3, dd=3
-Figure 1: Synthetic dataset and decision trees with corresponding 0-1 loss, tree sizes ($K$), and tree depths ($d$).
+Keywords: decision tree, dynamic programming, global optimal algorithm
+Figure 1: Synthetic dataset (left) generated by degree-2 polynomials; axis-parallel decision trees learned by CART and the optimal algorithm (middle two); and hypersurface decision trees learned using our proposed algorithm (right), with corresponding misclassification counts, tree sizes, and tree depths.
 1 Introduction
@@
 When the underlying decision regions involve complex or non-convex boundaries, trees with hyperplane or hypersurface splitting rules provide a more faithful representation, whereas axis-parallel trees remain limited to hyper-rectangular partitions.
-For instance, Figure LABEL:fig:_comparision shows a synthetic dataset generated from three quadratic hypersurfaces (conic sections).
+For instance, Figure 1 shows a synthetic dataset generated from three quadratic hypersurfaces (conic sections).
 The decision regions produced by classical CART and by the state-of-the-art optimal decision tree algorithm ConTree (Brița et al., 2025), displayed in the two middle panels, are constrained to irregular axis-aligned rectangles.
-Both CART and ConTree yield trees with 15 splitting rules, of depths 5 and 4, respectively.
+Both CART and ConTree yield trees with 15 splitting rules, of depths 5 and 4, respectively, whereas our method produce a tree with 3 splits of depth $3$.
 The above example motivates the design of algorithms for constructing decision trees with more flexible splitting rules, such as hyperplanes or, more generally, hypersurfaces.
 However, to the best of our knowledge, no algorithm—whether exact, approximate, or heuristic—has been proposed for learning decision trees with hypersurface splitting rules.
-In this second paper of the series, we present the first such algorithm, building on the results from Part I to solve the size-constrained optimal hypersurface decision tree (HODT) problem.
+In this second paper of the series, we present the first such algorithm, building on the results from Part I (He, 2025) to solve the size-constrained optimal hypersurface decision tree (HODT) problem.
 The algorithm is developed through several key results, including the prefix-closed properties of crossed hyperplanes and the incremental generation of ancestry relation matrices.
@@
 Part II, presented here, develops the first algorithm for solving the HODT problem, building directly on the foundations of Part I.
-Section 3, details the construction of the algorithm: Subsection 3.1 introduces ancestry relations for hyperplanes, identifying a critical property—crossed hyperplanes—and proving that any set of hyperplanes containing such a pair cannot form a proper decision tree.
+Section 3 details the construction of the algorithm: Subsection 3.1 introduces ancestry relations for hyperplanes, identifying a critical property—crossed hyperplanes—and proving that any set of hyperplanes containing such a pair cannot form a proper decision tree.
 This result forms the basis for an efficient ancestry-matrix generator free of crossed hyperplanes, described in Subsection 3.2.
@@
 General framework for solving the hypersurface decision tree problem
-In Part I, we formalized the decision tree problem through a set of axioms, referred to as the proper decision tree axiom.
+In Part I (He, 2025), we formalized the decision tree problem through a set of axioms, referred to as the proper decision tree axiom.
 Based on these axioms, we derived four generic definitions of the decision tree problem, expressed unambiguously as four recursive programs.
@@
 $\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\left[\;\right]\right)=\left[\mathit{DL}\left(\mathit{xs}\right)\right]$
-$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right]$
+$\mathit{genDTs}_{\text{vec}}$ $\left(\mathit{xs},\mathit{rs}\right)=\mathit{concat}\circ\left[\mathit{updates}\left(r,\mathit{rs}^{\prime},xs\right)|\left(r,\mathit{rs}^{\prime}\right)\longleftarrow\mathit{candidates}\left(\mathit{rs}\right)\right].$
 The definition in 3 is a two-phase process: $\mathit{min}_{E}$ is applied to $\mathit{genDTs}_{\text{vec}}$ only after all configurations are exhaustively generated.
@@
 In recalling the general framework for solving the size-constrained ODT problem, an astute reader may notice that the program $\mathit{odt}_{\text{size}}$ operates on a list of splitting rules $\mathit{rs}:\left[\mathcal{R}\right]$ rather than directly on the data.
-This design choice reflects our aim of providing a generic, simple, and modular framework, since decision trees are defined in terms of splitting rules rather than data instances.
+This design choice reflects our aim of providing a generic and modular framework, since decision trees are defined in terms of splitting rules rather than data instances.
 Consequently, when applying the framework to practical ODT problems in machine learning, a separate procedure $\mathit{gen}_{\mathit{splits}}$ must be introduced to generate the splitting rules, with its definition depending on the specific task.
@@
 $=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}\circ\mathit{embed}_{M}$
-$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M}$
+$=\mathit{min}_{E_{\text{0-1}}}\circ\mathit{concatMap}_{\mathit{sodt}\left(\mathit{xs}\right)}\circ\mathit{nestedCombs}\left(K,G\right)\circ\mathit{embed}_{M},$
 where $\mathit{embed}_{M}:\left[\mathbb{R}^{D}\right]\to\left[\mathbb{R}^{G}\right]$ is the Veronese embedding ($\rho_{M}$) introduced in Part I, mapping data points $\mathit{xs}$ in $\mathbb{R}^{D}$ to the embedded dataset $\rho_{M}\left(\mathit{xs}\right)$ in $\mathbb{R}^{G}$, with $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$.
-Moreover, $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ can be unified into a single program $\mathit{nestedCombs}\left(K,G\right)$, for which the implementation was given in Part I.
-For completeness, we include its pseudocode in Algorithm 5 in Appendix A.
+The transformation from the second to the final equation unifies $\mathit{kcombs}\left(K\right)\circ\mathit{gen}_{\mathit{splits}}^{M}$ into a single program $\mathit{nestedCombs}\left(K,G\right)$ as derived in Part I.
+For completeness, the pseudo-code for $\mathit{nestedCombs}\left(K,G\right)$ is provided in Algorithm 5.
 However, the specification (5) is not directly executable.
@@
 Thus, for any pair of axis-parallel hyperplanes $h_{i}^{0}$ and $h_{j}^{0}$ either $h_{i}^{0}\swarrow h_{j}^{0}$ or $h_{i}^{0}\searrow h_{j}^{0}$ holds.
-In other words, any axis-parallel hyperplane $h_{j}^{0}$ must lie in either the left or right subregion determined by another axis-parallel hyperplane $h_{i}^{0}$.
+In other words, any axis-parallel hyperplane $h_{j}^{0}$ must lie in either the left or right sub-region determined by another axis-parallel hyperplane $h_{i}^{0}$.
 In contrast, for general hyperplanes characterized by data points in general position, ancestry relations are more nuanced.
 The following fact highlights the key observation about the possible ancestry relations between any pair of such hyperplanes.
-Fact 1.
+Fact 1
 Let $h_{i}$ and $h_{j}$ be two hyperplanes defined by $D$ data points in general position. Their ancestry relation can fall into one of three categories:
@@
 To formalize this intuition, we define the concept of crossed hyperplane formally.
-Definition 1.
+Definition 1
 Crossed hyperplane.
@@
 Consequently, any combination of hyperplanes containing a crossed pair cannot yield a proper decision tree.
-Theorem 1.
+Theorem 2
 If two hyperplanes $h_{i}$ and $h_{j}$ cross each other then: no ancestry relation exists between $h_{i}$ and $h_{j}$, and no hyperplanes $h_{k}$ can separate $h_{i}$ and $h_{j}$ into different branches.
 Consequently, any combination of hypersurfaces containing such crossed hypersurfaces cannot form a proper decision tree.
-Proof.
-We prove this by contradiction.
+Proof We prove this by contradiction.
 Assume there exists a third hypersurface $h_{k}$ such that $h_{i}$ and $h_{j}$ lie on the left and right branches of $h_{k}$, respectively.
@@
 Hence, no hypersurface can separate crossed hypersurfaces, and any combination containing them cannot form a proper decision tree.
-∎
-Figure 4: An example illustrating the proof of Fact 1.
+Figure 4: An example illustrating the proof of Fact 2.
 Demonstrating that the data items $a$, $b$, which define $h_{i}$, and $c$, $d$, which define $h_{j}$ cannot be classified into the disjoint regions defined by a third hyperplane (red).
@@
 Any third hyperplane in $\mathbb{R}^{D}$, such as the red line in the figure, cannot separate the points $a,b$ and $c,d$ into distinct regions.
-Lemma 1.
+Lemma 3
 Given a list of $K$ hyperplanes $\mathit{hs}_{K}=\left[h_{1},h_{2},\ldots,h_{K}\right]$ with ancestry matrix $\boldsymbol{K}$.
@@
 If $\mathit{hs}_{K}$ contains no crossed hyperplanes, it can form at least one proper decision tree.
-Proof.
-By definition, if a pair $h_{i}$ and $h_{j}$ in $\mathit{hs}_{K}$ are crossed, then neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ nor $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable, and therefore $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$.
+Proof By definition, if a pair $h_{i}$ and $h_{j}$ in $\mathit{hs}_{K}$ are crossed, then neither $h_{i}\left(\swarrow\vee\searrow\right)h_{j}$ nor $h_{j}\left(\swarrow\vee\searrow\right)h_{i}$ is viable, and therefore $\boldsymbol{K}_{ij}=0$ and $\boldsymbol{K}_{ji}=0$.
 Conversely, if $\mathit{hs}_{K}$ contains no crossed hyperplanes, then for any pair $h_{i},h_{j}\in\mathit{hs}_{K}$, either a mutual ancestry or an asymmetrical ancestry exists.
 Hence, a proper decision tree can be constructed.
-∎
 A combination of hyperplanes $\mathit{hs}_{K}$ without crossed hyperplanes is referred to as a feasible combination.
@@
 This yields the following incremental (sequentially recursive) program $\mathit{calARM}$ for constructing $\mathit{ncr}_{K}$
-[…]
+$\mathit{calARM}$ $\left(\left[r\right]\right)=\left[\begin{array}[]{c}r\\ 0\end{array}\right]$
 $\mathit{calARM}$ $\left(r:\mathit{rs}\right)=\mathit{update}_{\text{arMat}}\left(r,\mathit{calARM}\left(\mathit{rs}\right)\right)$
@@
 Specifically, when the generator is defined sequentially, the following filter fusion theorem applies.
-Theorem 2.
+Theorem 4
 Filter fusion theorem. Let a sequential generator $\mathit{gen}$ be defined recursively as:
@@
 $p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$ (11)
-Proof.
-We prove the fusion theorem by following reasoning
+Proof We prove the fusion theorem by following reasoning
 $\mathit{filtgen}_{q}$
@@
 A configuration $a:\mathit{as}$ survived in $\mathit{filter}_{p}$ will also survived in $\mathit{filter}_{q}\left(\mathit{alg}_{2}\left(a,\mathit{filter}_{p}\left(\left[\mathit{as}\right]\right)\right)\right)$ because $p\left(a:\mathit{as}\right)=q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$.
-∎
 The reason the prefix-closed property (11) introduces an additional predicate $q$ is that, if we already know $p\left(\mathit{as}\right)$ holds, it is often more efficient to evaluate $q\left(a:\mathit{as}\right)\wedge p\left(\mathit{as}\right)$ rather than directly computing $p\left(a:\mathit{as}\right)$.
@@
 Similarly, for our problem, since $\mathit{nestedCombs}\left(K,D\right)$ can be defined sequentially, we can establish a prefix-closed property for the ancestry relation matrix, analogous to the auxiliary check in the eight queens problem.
-Fact 2.
+Fact 2
 Prefix-closed property for crossed hyperplanes.
@@
 This reduces the complexity to $O\left(K\times t\right)$.
-  1. 1.
-    Input: $r_{j}$: the index of the new splitting rule; $\mathit{ncsr}$: a $M\times\left(k+1\right)\times k$ tensor, which contains a vector of $\mathit{ncr}:NCR$ configurations; $\mathit{css}$: a matrix for storing all $G$-combinations of data points of all splitting rules; $\mathit{asgn}^{\pm}$: the positive/negative predictions of each splitting rule
-  2. 2.
-    Output: The updated nested combinations $\mathit{ncsr}^{\prime}:\mathit{NCRs}$ without crossed hyperplanes, which is a a vector of $\left(k+2\right)\times\left(k+1\right)$ matrices
-  3. 3.
-    $\mathit{ncrs}^{\prime}=\left[\>\right]$
-  4. 4.
-    for $\mathit{ncr}\in\mathit{ncrs}$ do:
-  5. 5.
-    for $r_{i}\in\mathit{ncr}\left[0\right]$ do:
-  6. 6.
-       $p_{1}=\text{True }\text{if }\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{+}\left[r_{i}\right]\vee\mathit{css}\left[G\right]\left[r_{j}\right]\in\mathit{asgn}^{-}\left[r_{i}\right]$ else False // the $G$-combination of data points $\mathit{css}\left[G\right]\left[r_{j}\right]$ all lies in the positive side or negative side of the hyperplane $r_{i}$,
-  7. 7.
-      $p_{2}=\text{True }\text{if }$ $\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{+}\left[r_{j}\right]\vee\mathit{unrank}\left(r_{i}\right)\in\mathit{asgn}^{-}\left[r_{j}\right]$ else False // the $G$-combination corresponds for defining rule $r_{i}$ all lies in the positive side or negative side of the splitting rule $r_{j}$, $\mathit{unrank}\left(r_{i}\right)$ takes an index and return all data points for defining splitting rule $r_{i}$,
-  8. 8.
-      if $p_{1}\vee p_{2}$ do:
-  9. 9.
-       $\mathit{ncr}^{\prime}=\mathit{update}_{\text{arMat}}\left(\mathit{ncr}\right)$ // update ancestry relation matrix
-  10. 10.
-       $\mathit{ncrs}^{\prime}+\!\!+\left[\mathit{ncr}^{\prime}\right]$
-  11. 11.
-      else:
-  12. 12.
-        break // break the inner loop because $\mathit{ncr}$ is infeasible
-  13. 13.
-    return $\mathit{ncrs}^{\prime}$
-Algorithm 1 $\mathit{updates}_{\text{arMat}}$
+Algorithm 1 Update Ancestry Relation ($\mathit{updates}_{arMat}$)
+1:Input: $\mathit{r}_{j}$: index of new splitting rule; $\mathit{ncrs}$: $M\times(k+1)\times k$ tensor of $\mathit{ncr}$ configurations; $\mathit{css}$: matrix of $G$-combinations; $\mathit{asgn}^{\pm}$: positive/negative predictions
+2:Output: Updated nested combinations $\mathit{ncrs}^{\prime}$: $\mathit{NCRs}$ without crossed hyperplanes, as $(k+2)\times(k+1)$ matrices
+[…]
+4:for $\mathit{ncr}\in\mathit{ncrs}$ do
+5:  for $\mathit{r}_{i}\in\mathit{ncr}[0]$ do
+6:   $p_{1}\leftarrow\top$ if $\mathit{css}[G][\mathit{r}_{j}]\in\mathit{asgn}^{+}[\mathit{r}_{i}]\vee\mathit{css}[G][\mathit{r}_{j}]\in\mathit{asgn}^{-}[\mathit{r}_{i}]$ else $\bot$ $\triangleright$ Check if $G$-combination lies on one side of $\mathit{r}_{i}$
+7:   $p_{2}\leftarrow\top$ if $\mathit{unrank}(\mathit{r}_{i})\in\mathit{asgn}^{+}[\mathit{r}_{j}]\vee\mathit{unrank}(\mathit{r}_{i})\in\mathit{asgn}^{-}[\mathit{r}_{j}]$ else $\bot$ $\triangleright$ Check if rule $\mathit{r}_{i}$ lies on one side of $\mathit{r}_{j}$
+8:   if $p_{1}\vee p_{2}$ then
+9:    $\mathit{ncr}^{\prime}\leftarrow\mathit{update}_{arMat}(\mathit{ncr})$ $\triangleright$ Update ancestry relation matrix
+10:    $\mathit{ncrs}^{\prime}\leftarrow\mathit{ncrs}^{\prime}+[\mathit{ncr}^{\prime}]$ $\triangleright$ Append updated configuration
+11:   else
+12:    break $\triangleright$ Skip infeasible $\mathit{ncr}$
+13:   end if
+14:  end for
+15:end for
+16:return $\mathit{ncrs}^{\prime}$
 We can now incorporate $\mathit{update}_{\text{arMat}}$ along with the additional filtering process for eliminating crossed hyperplanes directly into the definition of $\mathit{nestedCombs}\left(K,G\right)$.
@@
 This is achieved by introducing a single additional line of code after line 12 of Algorithm using Algorithm 1 to update all feasible nested combinations while avoiding non-feasible combinations containing crossed hyperplanes.
-The resulting modification is minimal—just one line—and the complete pseudocode is provided in Appendix A.2, Algorithm 6.
+The resulting modification is minimal—just one line—and the complete pseudo-code is provided in Appendix A.2, Algorithm 6.
 3.3 Optimal hypersurface decision tree algorithm
-  1. 1.
-    Input: $\mathit{xs}$: input data list of length $N$; $K$: number of splitting rules; $G$: dimension of the embedded space
-  2. 2.
-    Output: Array of $\left(k,D\right)$-nested-combinations for $k\in\left\{0,\ldots K\right\}$
-  3. 3.
-    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
-  4. 4.
-    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
-  5. 5.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
-  6. 6.
-    $\rho_{M}\left(\mathit{xs}\right)=\mathit{embed}\left(M,\mathit{xs}\right)$ // map the data list $\mathit{xs}$ to the embedded space
-  7. 7.
-    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
-  8. 8.
-    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
-  9. 9.
-      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
-  10. 10.
-      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
-  11. 11.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
-  12. 12.
-    $\mathit{css}\left[G\right]=\left[\>\right]$
-  13. 13.
-    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
-  14. 14.
-    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
-  15. 15.
-       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
-  16. 16.
-        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$
-  17. 17.
+Algorithm 2 OHDT with Thinning ($hodt(K,M,xs)$)
+1:Input: $\mathit{xs}$: data list, length NN; $K$: splitting rules; $G$: embedded space dimension
+2:Output: Array of $(k,D)$-nested-combinations for $k\in\{0,\ldots,K\}$
 […]
-  18. 18.
-    if $E_{\text{0-1}}\left(t_{\text{best}}\right)\leq E_{\text{0-1}}\left(t_{\text{opt}}\right)$:
-  19. 19.
-      $t_{\text{opt}}=t_{\text{best}}$
-  20. 20.
-    $\mathit{ncss}\left[K\right]=\left[\>\right]$ // eliminate size $K$ nested combinations after evaluation
-  21. 21.
-    return $t_{\text{opt}}$
-Algorithm 2 $\mathit{hodt}\left(K,M,\mathit{xs}\right)$
+[…]
+[…]
+[…]
+7:for $n\in\mathit{range}(0,N)$ do
+8:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
+9:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
+10:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
+11:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
+12:   $\mathit{css}[G]\leftarrow[\,]$
+13:  end for
+14:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n}{G-1}$
+15:  for $i\in\mathit{range}(C_{1},C_{2})$ do
+16:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
+17:    $\mathit{ncss}[k]\leftarrow\mathit{updates}_{arMat}(i,\mathit{ncss}[k],\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$
+18:   end for
+19:  end for
+20:  $t_{best}\leftarrow\min_{E_{0-1}}(\mathit{mapL}(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}[K]))$ $\triangleright$ Optimal decision tree
+21:  if $E_{0-1}(t_{best})\leq E_{0-1}(t_{opt})$ then
+22:   $t_{opt}\leftarrow t_{best}$
+23:  end if
+24:  $\mathit{ncss}[K]\leftarrow[\,]$ $\triangleright$ Clear size $K$ nested combinations
+25:end for
+26:return $t_{opt}$
 Finally, after introducing the prefix-closed filtering and incremental update for ancestry relation matrix, we are now ready to fuse all components into a single program.
@@
 In the specification, functions are composed sequentially, meaning each function begins only after the previous one has completed.
-If written as pseudocode, both $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ would appear outside the outermost loop that defines $\mathit{\mathit{nestedCombs}}\left(K,G\right)$.
+If written as pseudo-code, both $\mathit{min}{}_{E_{\text{0-1}}}$ and $\mathit{concatMap}{}_{sodt\left(\mathit{xs}\right)}$ would appear outside the outermost loop that defines $\mathit{\mathit{nestedCombs}}\left(K,G\right)$.
 This is inefficient because $\mathit{\mathit{nestedCombs}}\left(K,G\right)$ generates a prohibitively large number of candidate solutions.
@@
 We hope that the rigor invested in designing this optimal algorithm will positively influence future studies, demonstrating not only how formal steps can guarantee optimality but also the advantages of adopting a formal, principled approach.
-4 Heuerestic methods
+4 Heuristic methods
 The use of heuristics is common in the study of optimal algorithms.
@@
 In this section, we develop two simple yet effective heuristics for addressing the HODT problem.
-  1. 1.
-    Input: Parameteres for coreset: $\mathit{BS}$: Block size; $R$: number of shuffle time in each filtering process; $L$: Max-heap size; $B_{\max}$: Maximum input size for the Deep-ICE algorithm; $c\in\left(0,1\right]$: Shrinking factor for heap size
-  2. 2.
-      Parameteres for $\mathit{hodt}$: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\mathit{xs}$: input data list;
-  3. 3.
-    Output: Max-heap containing top $L$ configurations and associated data blocks
-  4. 4.
-    Initialize coreset $\mathcal{C}\leftarrow\mathit{ds}$
-  5. 5.
-    while $\mathcal{C}\leq B_{\max}$ do:
-  6. 6.
-    Reshuffle the data, divide $\mathcal{C}$ into $\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil$ blocks $\mathcal{C}_{B}=\left\{C_{1},C_{2},\ldots,C_{\left\lceil\frac{\left|\mathcal{C}\right|}{\mathit{BS}}\right\rceil}\right\}$
-  7. 7.
-    Initialize a size $L$ max-heap $\mathcal{H}_{L}$
-  8. 8.
-    for $r\leftarrow 1$ to $R$ do:
-  9. 9.
-      $r=r+1$
-  10. 10.
-      for $C\in\mathcal{C}_{B}$ do:
-  11. 11.
-       $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,C\right)$
-  12. 12.
-       $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},C\right)$
-  13. 13.
-    $\mathcal{C}\leftarrow\mathit{unique}\left(\mathcal{H}_{L}\right)$ // Merge blocks and remove duplicates
-  14. 14.
-    $L\leftarrow L\times c$ // Shrink heap size:
-  15. 15.
-    $\mathit{cnfg}\leftarrow\mathit{hodt}\left(K,M,\mathcal{C}\right)$ // Final refinement
-  16. 16.
-    $\mathcal{H}_{L}.\text{push}\left(\mathit{cnfg},\mathcal{C}\right)$
-  17. 17.
-    return $\mathcal{H}_{L}$
-Algorithm 3 $\mathit{hodt}Coreset\left(K,M,\mathit{xs},\mathit{BS},R,L,B_{\max},c\right)$
 4.1 Coreset selection method
+Algorithm 3 Coreset Selection ($\mathit{hodtCoreset}(K,M,\mathit{xs},\mathit{BS},R,L,B_{\max},c)$)
+1:Input: Coreset parameters: $\mathit{BS}$: block size; $R$: shuffle iterations; $L$: max-heap size; $B_{\max}$: max input size; $c\in(0,1]$: shrinking factor
+2:      $\mathit{hodt}$ parameters: $K$: splitting rules; $M$: polynomial degree; $\mathit{xs}$: data list
+3:Output: Max-heap with top $L$ configurations and data blocks
+[…]
+5:while $|\mathcal{C}|\leq B_{\max}$ do
+6:  $\mathcal{C}_{B}\leftarrow\{C_{1},C_{2},\ldots,C_{\lceil|\mathcal{C}|/\mathit{BS}\rceil}\}$ $\triangleright$ Reshuffle and divide into blocks
+7:  $\mathcal{H}_{L}\leftarrow\emptyset_{L}$ $\triangleright$ Initialize size $L$ max-heap
+8:  for $r\in\mathit{range}(1,R)$ do
+9:   for $C\in\mathcal{C}_{B}$ do
+10:    $\mathit{cnfg}\leftarrow\mathit{hodt}(K,M,C)$ $\triangleright$ Compute configuration
+11:    $\mathcal{H}_{L}.\mathit{push}(\mathit{cnfg},C)$ $\triangleright$ Push to heap
+12:   end for
+13:  end for
+14:  $\mathcal{C}\leftarrow\mathit{unique}(\mathcal{H}_{L})$ $\triangleright$ Merge blocks, remove duplicates
+15:  $L\leftarrow L\times c$ $\triangleright$ Shrink heap size
+16:end while
+[…]
+[…]
+19:return $\mathcal{H}_{L}$
 The first method, as reported by He et al.
@@
 4.2 $\mathit{sodt}$ with selected hyperplanes
-  1. 1.
-    Input: $K$: number of splitting rules, $M$ degree of polynomial hypersurface splitting rules; $\alpha$: threshold for duplicate data $\mathit{xs}$: input data list;
-  2. 2.
-    Output: The updated nested combinations $\mathit{NCs}$ without crossed hyperplanes, a vector of e $\left(k+2\right)\times\left(k+1\right)$ matrices
-  3. 3.
-    $t_{\text{best}}=\mathit{empty}\left(K,1\right)$ // a vector
-  4. 4.
-    $\mathcal{H}_{C}=\mathit{\mathit{hodt}Coreset}\left(1,M,\mathit{xs},BS,R,L,B_{\max},c\right)$ // selecting candidate hypersurfaces with low 0-1 loss
-  5. 5.
-    $G=\left(\begin{array}[]{c}M+D\\ D\end{array}\right)-1$
-  6. 6.
-    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
-  7. 7.
-    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
-  8. 8.
-    for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,n+1\right)\right)$ do:
-  9. 9.
-      $\mathit{ncss}^{\prime}=\mathit{filter}_{q\left(\alpha\right)}\left(\mathit{map}\left(\cup\left[n\right],\mathit{ncss}\left[k-1\right]\right)\right)$ // filter out nested combinations $\mathit{nc}$ in $\mathit{ncss}^{\prime}\left[k\right]$ with duplicates data points small than threshold $\alpha$
-  10. 10.
-      $t_{\text{best}}\left[k\right]=\mathit{min}_{E_{\text{0-1}}}\left(\mathit{mapL}\left(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}^{\prime}\right)\right)$
-  11. 11.
-      $\mathit{ncss}\left[k\right]=\mathit{ncss}\left[k\right]\cup\mathit{reverse}\left(\mathit{ncss}^{\prime}\right)$ // update $\mathit{ncss}$ to generate combinations in revolving door ordering,
-  12. 12.
-    $\mathit{ncss}\left[K\right]=\left[\>\right]$
-  13. 13.
-    return $t_{\text{best}}$
-Algorithm 4 $\mathit{sodtWSH}\left(K,M,\alpha,\mathit{xs}\right)$
+Algorithm 4 $\mathit{sodt}$ with selected hyperplanes ($\mathit{sodtWSH}(K,M,\alpha,\mathit{xs})$)
+1:Input: $K$: splitting rules; $M$: polynomial degree; $\alpha$: duplicate threshold; $\mathit{xs}$: data list
+2:Output: Nested combinations $\mathit{NCs}$ without crossed hyperplanes, as $(k+2)\times(k+1)$ matrices
+[…]
+[…]
+[…]
+[…]
+7:for $n\in\mathit{range}(0,N)$ do
+8:  for $k\in\mathit{reverse}(\mathit{range}(K,n+1))$ do
+9:   $\mathit{ncss}^{\prime}\leftarrow\mathit{filter}_{q(\alpha)}(\mathit{map}(\cup[n],\mathit{ncss}[k-1]))$ $\triangleright$ Filter duplicates below $\alpha$
+10:   $t_{best}[k]\leftarrow\min_{E_{0-1}}(\mathit{mapL}(\mathit{sodt}_{\mathit{xs}},\mathit{ncss}^{\prime}))$ $\triangleright$ Compute best configuration
+11:   $\mathit{ncss}[k]\leftarrow\mathit{ncss}[k]\cup\mathit{reverse}(\mathit{ncss}^{\prime})$ $\triangleright$ Update combinations
+12:   $\mathit{ncss}[K]\leftarrow[\,]$ $\triangleright$ Clear size $K$ combinations
+13:  end for
+14:end for
+15:return $t_{best}$
 Although the coreset selection method provides plausible solutions for low-dimensional datasets, an obvious limitation arises for high-dimensional datasets: the number of possible splitting rules is too large, even when the dataset is partitioned into small blocks.
@@
 In this section, we analyze in detail the computational efficiency of $\mathit{sodt}$ for solving the size-constrained ODT problem.
-Since $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ share similar advantages—both are fully vectorizable but unable to exploit efficient dynamic programming (DP)—we focus on comparing $\mathit{sodt}_{\text{vec}}$ against the dynamic programming algorithm $\mathit{sodt}_{\text{rec}}$.
+Since $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{kperms}}$ share similar advantages—both are fully vectorized but unable to exploit efficient dynamic programming (DP)—we focus on comparing $\mathit{sodt}_{\text{vec}}$ against the dynamic programming algorithm $\mathit{sodt}_{\text{rec}}$.
 Rather than using a binary tree data structure, which relies on pointers to locate subtrees and suffers from poor cache performance, both $\mathit{sodt}_{\text{vec}}$ and $\mathit{sodt}_{\text{rec}}$ are implemented entirely with array (heap) data structures.
@@
 5.1.1 Comparison between the vectorized and recursive implementation
-Figure 5: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on sequential setting. .
-Figure 6: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on parallel setting..
+Figure 5: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on sequential setting.
+Figure 6: Running time comparison between $\mathit{sodt}_{\text{rec}}$ and $\mathit{sodt}_{\text{vec}}$ with varying $K$ on parallel setting.
 Even though $\mathit{sodt}_{\text{rec}}$ has provably lower theoretical complexity, the hardware compatibility of $\mathit{sodt}_{\text{vec}}$ allows it to outperform in practice, particularly as $K$ increases.
@@
 We focus on analyzing the pure computational scalability of the algorithm by benchmarking its performance without any acceleration techniques.
-Specifically, we measure the total wall-clock runtime of $\mathit{sodt}_{\text{vec}}$ for solving $1\times 10^{6}$ instances of feasible nested combinations (with no crossed hyperplanes) for different values of KK。In our setup, cases with $K\leq 4$ are executed on the CPU, while cases with $K>4$ are executed on the GPU.
+Specifically, we measure the total wall-clock runtime of $\mathit{sodt}_{\text{vec}}$ for solving $1\times 10^{6}$ instances of feasible nested combinations (with no crossed hyperplanes) for different values of $K$.
+In our setup, cases with $K\leq 4$ are executed on the CPU, while cases with $K>4$ are executed on the GPU.
 As expected, when $K$ is fixed $\mathit{sodt}_{\text{vec}}$ (with worst-case complexity $O\left(K!\times N\right)$) exhibits linear runtime growth, which appears logarithmic in a log-linear plot.
@@
 5.2 Combinatorial complexity of the hyperplane/hypersurface decision tree
-5.2.1 Combinatorial complexity of nested combinations after fitlered out crossed hyperplanes
+5.2.1 Combinatorial complexity of nested combinations after filtered out crossed hyperplanes
 Figure 8: Combinatorial complexity of feasible nested combinations with varying $K$.
-Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
+Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) to the value of $K$ labeled below each group of bars.
 As noted, the combinatorial complexity of the hyperplane ($M=1$) ODT problem is bounded by $\left(\begin{array}[]{c}\left(\begin{array}[]{c}N\\ D\end{array}\right)\\ K\end{array}\right)=O\left(N^{DK}\right)$.
@@
 Nevertheless, in practice, this behavior is rarely observed, as it typically requires $K$ to reach astronomically large values even for medium-sized datasets.
-Comparing the combinatorial complexity of hypersurface and hyperplanes in the same diemension
+Comparing the combinatorial complexity of hypersurface and hyperplanes in the same dimension
 Figure 9: Combinatorial complexity of feasible nested combinations with varying $K$.
-Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) (leftmost bar) to the value of $K$ labeled below each group of bars.
+Each bar corresponds to a fixed $K$ starts from $K=2$ (left-most bar) to the value of $K$ labeled below each group of bars.
 Having analyzed the combinatorial complexity of hyperplane splitting rules, we now extend the analysis to hypersurface rules by comparing their complexity with that of hyperplanes in the same dimension.
@@
 Maximum Average Expected
-2
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.12 (0.04)
-93.82 (3.01)
-91.72 (3.83)
-100 (0.00)
-100 (0.00)
-82.61 (0.04)
-86.12 (3.73)
-84.57 (4.03)
-93.50 (3.21)
-100 (0.00)
-2.92 (0.27)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.18 (0.38)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.97 (0.09)
-2.49 (0.16)
-2.00 (0.00)
-1.67 (1.20)
-1.67 (1.20)
-1.96 (0.15)
-2.15 (0.20)
-2.00 (0.00)
-1.65 (0.16)
-1.66 (0.16)
-3
-CART-depth-3
-CART-size-6
-ConTree-3
-HODT
-Ground Truth
-90.72 (047)
-93.88 (3.89)
-95.98 (2.93)
-100 (0.00)
-100 (0.00)
-81.30 (5.71)
-83.92 (5.55)
-85.25 (4.96)
-94.91 (2.36)
-1 (0.00)
-5.90 (0.00)
-6.00 (0.00)
-6.94 (0.24)
-3.00 (0.00)
-3.00 (0.00)
-3.00 (0.00)
-4.02 (0.65)
-3.00 (0.00)
-3.00 (0.00)
-2.72 (0.45)
-2.83 (0.14)
-3.10 (0.24)
-2.99 (0.03)
-2.25 (0.0)
-2.18 (0.11)
-2.77 (0.19)
-2.71 (0.20)
-2.99 (0.04)
-2.18 (0.25)
-2.07 (0.24)
-4
-CART-depth-3
-CART-size-8
-ConTree-3
-HODT
-Ground Truth
-87.98 (5.26)
-93.90 (3.93)
-93.12 (4.02)
-97.46 (1.70)
-100 (0.00)
-77.66 (5.38)
-81.93 (4.93)
-81.51 (4.78)
-89.89 (3.71)
-100 (0.00)
-6.12 (90.86)
-8.00 (0.00)
-7.00 (0.00)
-4.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-4.54 (0.57)
-3.00 (0.00)
-3.52 (0.50)
-3.52 (0.50)
-2.87 (0.14)
-3.49 (0.21)
-3.00 (0.00)
-2.63 (0.19)
-2.63 (0.19)
-2.86 (0.15)
-3.13 (0.21)
-3.00 (0.00)
-2.60 (0.31)
-2.63 (0.33)
+2 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.12 (0.04) 93.82 (3.01) 91.72 (3.83) 100 (0.00) 100 (0.00) 82.61 (0.04) 86.12 (3.73) 84.57 (4.03) 93.50 (3.21) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.49 (0.16) 2.00 (0.00) 1.67 (1.20) 1.67 (1.20) 1.96 (0.15) 2.15 (0.20) 2.00 (0.00) 1.65 (0.16) 1.66 (0.16)
+3 CART-depth-3 CART-size-6 ConTree-3 HODT Ground Truth 90.72 (047) 93.88 (3.89) 95.98 (2.93) 100 (0.00) 100 (0.00) 81.30 (5.71) 83.92 (5.55) 85.25 (4.96) 94.91 (2.36) 100 (0.00) 5.90 (0.00) 6.00 (0.00) 6.94 (0.24) 3.00 (0.00) 3.00 (0.00) 3.00 (0.00) 4.02 (0.65) 3.00 (0.00) 3.00 (0.00) 2.72 (0.45) 2.83 (0.14) 3.10 (0.24) 2.99 (0.03) 2.25 (0.00) 2.18 (0.11) 2.77 (0.19) 2.71 (0.20) 2.99 (0.04) 2.18 (0.25) 2.07 (0.24)
+4 CART-depth-3 CART-size-8 ConTree-3 HODT Ground Truth 87.98 (5.26) 93.90 (3.93) 93.12 (4.02) 97.46 (1.70) 100 (0.00) 77.66 (5.38) 81.93 (4.93) 81.51 (4.78) 89.89 (3.71) 100 (0.00) 6.12 (90.86) 8.00 (0.00) 7.00 (0.00) 4.00 (0.00) 4.00 (0.00) 3.00 (0.00) 4.54 (0.57) 3.00 (0.00) 3.52 (0.50) 3.52 (0.50) 2.87 (0.14) 3.49 (0.21) 3.00 (0.00) 2.63 (0.19) 2.63 (0.19) 2.86 (0.15) 3.13 (0.21) 3.00 (0.00) 2.60 (0.31) 2.63 (0.33)
 Table 1: The effect of tree size of ground truth tree. No noise in data, training size = 100.
@@
 Maximum Average Expected
-100
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.76 (5.11)
-97.62 (2.51)
-93.52 (3.53)
-100 (0.00)
-100 (0.00)
-75.46 (15.70)
-80.26 (16.80)
-78.62 (16.14)
-84.38 (19.27)
-100 (0.00)
-2.88 (0.32)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-0.00 (0.00)
-2.00 (0.00)
-3.92 (0.69)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.96 (0.11)
-3.01 (0.30)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.96 (0.12)
-2.52 (0.33)
-2.00 (0.00)
-1.66 (0.25)
-1.70 (0.25)
-200
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.00 (4.49)
-94.49 (3.02)
-92.87 (3.38)
-99.94 (0.16)
-100 (0.00)
-84.62 (5.23)
-89.57 (4.04)
-88.51 (3.99)
-97.97 (1.01)
-100 (0.00)
-2.92.00 (0.27)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.26 (0.44)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.97 (0.09)
-2.50 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.98 (0.07)
-2.17 (0.21)
-2.00 (0.00)
-1.71 (0.25)
-1.70 (0.25)
-400
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-87.45 (4.69)
-92.58 (3.48)
-91.44 (3.37)
-99.85 (0.21)
-100 (0.00)
-84.62 (4.90)
-89.35 (4.08)
-88.78 (4.00)
-98.98 (0.58)
-100 (0.00)
-2.98 (0.14)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.24 (0.43)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.05)
-2.53 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.99 (0.05)
-2.15 (0.24)
-2.00 (0.00)
-1.71 (0.25)
-1.70 (0.25)
-800
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-86.91 (4.63)
-92.21 (3.43)
-91.00 (3.59)
-99.87 (0.14)
-100 (0.00)
-85.14 (4.96)
-90.12 (3.82)
-89.18 (3.94)
-99.40 (0.36)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.3 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.53 (0.18)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.13 (0.22)
-2.00 (0.00)
-1.69 (0.25)
-1.70 (0.25)
-1600
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-85.66 (5.05)
-91.48 (3.69)
-90.13 (3.69)
-99.87 (0.12)
-100 (0.00)
-84.97 (5.15)
-90.53 (3.71)
-89.38 (3.84)
-99.65 (0.22)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.18 (0.38)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.51 (0.16)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.16 (0.26)
-2.00 (0.00)
-1.70 (0.24)
-1.70 (0.25)
+100 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.76 (5.11) 97.62 (2.51) 93.52 (3.53) 100 (0.00) 100 (0.00) 75.46 (15.70) 80.26 (16.80) 78.62 (16.14) 84.38 (19.27) 100 (0.00) 2.88 (0.32) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 0.00 (0.00) 2.00 (0.00) 3.92 (0.69) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.96 (0.11) 3.01 (0.30) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.96 (0.12) 2.52 (0.33) 2.00 (0.00) 1.66 (0.25) 1.70 (0.25)
+200 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.00 (4.49) 94.49 (3.02) 92.87 (3.38) 99.94 (0.16) 100 (0.00) 84.62 (5.23) 89.57 (4.04) 88.51 (3.99) 97.97 (1.01) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.50 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.98 (0.07) 2.17 (0.21) 2.00 (0.00) 1.71 (0.25) 1.70 (0.25)
+400 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 87.45 (4.69) 92.58 (3.48) 91.44 (3.37) 99.85 (0.21) 100 (0.00) 84.62 (4.90) 89.35 (4.08) 88.78 (4.00) 98.98 (0.58) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.05) 2.15 (0.24) 2.00 (0.00) 1.71 (0.25) 1.70 (0.25)
+800 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 86.91 (4.63) 92.21 (3.43) 91.00 (3.59) 99.87 (0.14) 100 (0.00) 85.14 (4.96) 90.12 (3.82) 89.18 (3.94) 99.40 (0.36) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.3 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.53 (0.18) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.13 (0.22) 2.00 (0.00) 1.69 (0.25) 1.70 (0.25)
+1600 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 85.66 (5.05) 91.48 (3.69) 90.13 (3.69) 99.87 (0.12) 100 (0.00) 84.97 (5.15) 90.53 (3.71) 89.38 (3.84) 99.65 (0.22) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.51 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.16 (0.26) 2.00 (0.00) 1.70 (0.24) 1.70 (0.25)
 Table 2: The effect of training data size. Ground truth-tree size $K=2$.
@@
 Maximum Average Expected
-2
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.00 (4.49)
-94.49 (3.02)
-92.87 (3.38)
-99.94 (0.16)
-100 (0.00)
-84.62 (5.23)
-89.57 (4.04)
-88.51 (3.99)
-97.97 (1.01)
-100 (0.00)
-2.92 (0.27)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.26 (0.44)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.97 (0.09)
-2.53 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.98 (0.07)
-2.17 (0.21)
-2.0 (0.00)
-1.71 (0.25)
-1.70 (0.25)
-4
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-79.94 (4.53)
-84.17 (4.41)
-82.80 (4.16)
-95.13 (1.38)
-100 (0.00)
-73.97 (4.74)
-76.72 (4.48)
-76.16 (4.82)
-90.97 (2.55)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.10 (0.3)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.48 (0.13)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.24 (0.18)
-2.00 (0.00)
-1.67 (0.13)
-1.67 (0.12)
-8
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-75.05 (5.72)
-79.25 (5.66)
-77.78 (5.36)
-89.08 (4.26)
-100 (0.00)
-67.54 (7.08)
-69.66 (6.75)
-69.40 (6.37)
-82.95 (5.70)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.12 (0.325)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.48 (0.14)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.21 (0.22)
-2.00 (0.00)
-1.48 (0.12)
-1.49 (0.13)
-12
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-70.39 (3.82)
-74.50 (4.00)
-73.35 (3.49)
-85.06 (5.19)
-100 (0.00)
-62.78 (5.20)
-64.03 (4.99)
-63.53 (4.78)
-76.71 (6.47)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.14 (0.35)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.46 (0.14)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.26 (0.13)
-2.00 (0.00)
-1.50 (0.11)
-1.52 (0.10)
-Figure 10: Effects of dimensionality. Training size = 100. No noise. Ground truth tress are size 2.
+2 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.00 (4.49) 94.49 (3.02) 92.87 (3.38) 99.94 (0.16) 100 (0.00) 84.62 (5.23) 89.57 (4.04) 88.51 (3.99) 97.97 (1.01) 100 (0.00) 2.92 (0.27) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.98 (0.07) 2.17 (0.21) 2.0 (0.00) 1.71 (0.25) 1.70 (0.25)
+4 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 79.94 (4.53) 84.17 (4.41) 82.80 (4.16) 95.13 (1.38) 100 (0.00) 73.97 (4.74) 76.72 (4.48) 76.16 (4.82) 90.97 (2.55) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.10 (0.3) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.48 (0.13) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.24 (0.18) 2.00 (0.00) 1.67 (0.13) 1.67 (0.12)
+8 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 75.05 (5.72) 79.25 (5.66) 77.78 (5.36) 89.08 (4.26) 100 (0.00) 67.54 (7.08) 69.66 (6.75) 69.40 (6.37) 82.95 (5.70) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.12 (0.325) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.48 (0.14) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.21 (0.22) 2.00 (0.00) 1.48 (0.12) 1.49 (0.13)
+12 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 70.39 (3.82) 74.50 (4.00) 73.35 (3.49) 85.06 (5.19) 100 (0.00) 62.78 (5.20) 64.03 (4.99) 63.53 (4.78) 76.71 (6.47) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.14 (0.35) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.46 (0.14) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.26 (0.13) 2.00 (0.00) 1.50 (0.11) 1.52 (0.10)
+Table 3: Effects of dimensionality. Training size = 100. No noise. Ground truth tress are size 2.
 The third set of experiments examines the effect of problem dimensionality while keeping the tree size ($K=2$) and training set size ($N=100$) constant.
@@
 Effect of noise on labels
-Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
+Noise level Method Train Acc (%) Test Acc (%) Tree size Depth
 Maximum Average Expected
-0
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-90.42 (4.61)
-95.19 (3.11)
-93.49 (3.60)
-99.62 (0.68)
-1 (0.00)
-87.25 (5.20)
-91.21 (3.94)
-89.99 (3.71)
-98.06 (1.21)
-1 (0.00)
-2.92 (0.27)
-4.0 (0.00)
-2.98 (0.04)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.18 (0.38)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.97 (0.09)
-2.49 (0.16)
-1.99 (0.05)
-1.67 (0.00)
-1.67 (0.00)
-1.99 (0.03)
-2.11 (0.23)
-1.98 (0.11)
-1.67 (0.28)
-1.68 (0.28)
-5
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-86.29 (4.48)
-90.64 (3.07)
-89.07 (3.45)
-95.00 (0.20)
-0.95 (0.00)
-87.15 (4.89)
-91.01 (3.67)
-89.87 (3.47)
-97.20 (1.40)
-1 (0.00)
-2.94 (0.24)
-4.0 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.2.00 (0.40)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.98 (0.08)
-2.50 (0.16)
-2.00 (0.00)
-1.68 (0.00)
-1.67 (0.00)
-1.99 (0.02)
-2.10 (0.24)
-2.00 (0.00)
-1.62 (0.29)
-1.68 (0.28)
-10
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-82.25 (3.78)
-86.25 (2.67)
-84.69 (3.09)
-90.45 (0.49)
-0.90 (0.00)
-86.74 (5.36)
-90.45 (4.07)
-89.48 (3.96)
-97.59 (1.21)
-100 (0.00)
-2.98 (0.14)
-4.0 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.12 (0.32)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.05)
-3.12 (0.32)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.01)
-2.16 (0.26)
-2.00 (0.00)
-1.64 (0.30)
-1.68 (0.28)
-15
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-77.90 (3.99)
-81.80 (3.10)
-80.53 (3.04)
-85.69 (0.54)
-84.99 (0.00)
-85.99 (5.71)
-90.19 (3.67)
-89.10 (3.79)
-97.10 (1.84)
-100 (0.00)
-2.98 (0.14)
-4.0 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.22 (0.41)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.05)
-2.51 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.04)
-2.19 (0.31)
-2.00 (0.00)
-1.63 (0.29)
-1.68 (0.28)
-20
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-73.88 (3.16)
-77.36 (2.91)
-76.28 (2.81)
-80.98 (0.69)
-79.90 (0.00)
-86.06 (4.95)
-89.51 (4.59)
-88.61 (4.43)
-96.58 (2.06)
-100 (0.00)
-2.94 (0.24)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.24 (0.43)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.98 (0.08)
-2.54 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.01)
-2.13 (0.34)
-2.00 (0.00)
-1.68 (0.27)
-1.68 (0.28)
-25
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-70.35 (3.22)
-73.19 (2.54)
-72.26 (2.62)
-76.37 (0.83)
-75.00 (0.00)
-84.59 (6.66)
-87.78 (5.41)
-87.06 (5.16)
-95.67 (2.93)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.26 (0.44)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.52 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.24 (0.32)
-2.00 (0.00)
-1.67 (0.78)
-1.68 (0.28)
+0 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.42 (4.61) 95.19 (3.11) 93.49 (3.60) 99.62 (0.68) 100 (0.00) 87.25 (5.20) 91.21 (3.94) 89.99 (3.71) 98.06 (1.21) 100 (0.00) 2.92 (0.27) 4.0 (0.00) 2.98 (0.04) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.97 (0.09) 2.49 (0.16) 1.99 (0.05) 1.67 (0.00) 1.67 (0.00) 1.99 (0.03) 2.11 (0.23) 1.98 (0.11) 1.67 (0.28) 1.68 (0.28)
+5 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 86.29 (4.48) 90.64 (3.07) 89.07 (3.45) 95.00 (0.20) 95.00 (0.00) 87.15 (4.89) 91.01 (3.67) 89.87 (3.47) 97.20 (1.40) 100 (0.00) 2.94 (0.24) 4.0 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.2.00 (0.40) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.98 (0.08) 2.50 (0.16) 2.00 (0.00) 1.68 (0.00) 1.67 (0.00) 1.99 (0.02) 2.10 (0.24) 2.00 (0.00) 1.62 (0.29) 1.68 (0.28)
+10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 82.25 (3.78) 86.25 (2.67) 84.69 (3.09) 90.45 (0.49) 90.00 (0.00) 86.74 (5.36) 90.45 (4.07) 89.48 (3.96) 97.59 (1.21) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.12 (0.32) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 3.12 (0.32) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.16 (0.26) 2.00 (0.00) 1.64 (0.30) 1.68 (0.28)
+15 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 77.90 (3.99) 81.80 (3.10) 80.53 (3.04) 85.69 (0.54) 84.99 (0.00) 85.99 (5.71) 90.19 (3.67) 89.10 (3.79) 97.10 (1.84) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.22 (0.41) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.51 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.04) 2.19 (0.31) 2.00 (0.00) 1.63 (0.29) 1.68 (0.28)
+20 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 73.88 (3.16) 77.36 (2.91) 76.28 (2.81) 80.98 (0.69) 79.90 (0.00) 86.06 (4.95) 89.51 (4.59) 88.61 (4.43) 96.58 (2.06) 100 (0.00) 2.94 (0.24) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.98 (0.08) 2.54 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.13 (0.34) 2.00 (0.00) 1.68 (0.27) 1.68 (0.28)
+25 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 70.35 (3.22) 73.19 (2.54) 72.26 (2.62) 76.37 (0.83) 75.00 (0.00) 84.59 (6.66) 87.78 (5.41) 87.06 (5.16) 95.67 (2.93) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.26 (0.44) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.52 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.24 (0.32) 2.00 (0.00) 1.67 (0.78) 1.68 (0.28)
 Table 4: The effect of noise on labels. Training data size = 100. Ground truth trees are size 2.
@@
 From a broader perspective, when increasing noise from 0% to 25%, HODT’s out-of-sample accuracy decreased by 2.39%, compared with 2.93%, 3.43%, and 2.66% for ConTree, CART-size, and CART-depth, respectively.
-Notably, starting at a 10% noise level, HODT was able to find solutions that exceeded the ground truth on the training dataset while maintaining strong performance on out-of-sample tests.
+Notably, starting at a 5% noise level, HODT was able to find solutions that exceeded the ground truth on the training dataset while maintaining strong performance on out-of-sample tests.
 These results further refute the notion that optimal methods are less robust to noise, highlighting the superiority of HODT.
 Effect of noise on data
-Noise level (%) Method Train Acc (%) Test Acc (%) Tree size Depth
+Noise level Method Train Acc (%) Test Acc (%) Tree size Depth
 Maximum Average Expected
-0
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-90.29 (4.74)
-94.20 (3.70)
-92.73 (3.90)
-99.88 (0.26)
-100 (0.00)
-87.23 (5.14)
-90.32 (4.24)
-89.55 (4.10)
-98.08 (1.00)
-100 (0.00)
-2.96 (0.20)
-4.0 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.2 (0.399)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.065)
-2.51 (0.16)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.99 (0.09)
-2.09 (0.33)
-2.00 (0.00)
-1.63 (0.29)
-1.63 (0.28)
-5
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-90.16 (4.63)
-93.98 (3.78)
-92.62 (3.79)
-99.62 (0.40)
-99.59 (0.44)
-87.16 (5.06)
-90.32 (4.20)
-89.50 (3.94)
-98.10 (1.01)
-100 (0.00)
-2.98 (0.14)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.24 (0.43)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.05)
-2.53 (0.168)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.01)
-2.05 (0.32)
-2.00 (0.00)
-1.64 (0.28)
-1.65 (0.28)
-10
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-90.30 (4.55)
-93.98 (3.57)
-92.50 (3.65)
-99.07 (0.62)
-98.81 (0.71)
-87.36 (5.09)
-90.30 (4.55)
-89.38 (4.09)
-97.50 (0.90)
-100 (0.00)
-2.96 (0.20)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.28 (0.45)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.065)
-2.54 (0.17)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-1.99 (0.09)
-2.03 (0.34)
-2.00 (0.00)
-1.58 (0.29)
-1.65 (0.28)
-15
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.86 (4.91)
-93.78 (3.92)
-92.52 (3.83)
-99.05 (0.66)
-98.72 (0.80)
-87.06 (4.97)
-90.27 (4.42)
-89.31 (4.38)
-97.90 (1.10)
-100 (0.00)
-2.96 (0.20)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.32 (0.47)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.07)
-2.55 (0.18)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.02)
-2.03 (0.33)
-2.00 (0.00)
-1.63 (0.28)
-1.65 (0.28)
-20
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.75 (4.98)
-93.75 (3.70)
-92.39 (3.72)
-98.65 (0.87)
-98.11 (1.04)
-86.61 (5.57)
-90.21 (4.27)
-89.25 (4.05)
-97.40 (11.50)
-100 (0.00)
-3.00 (0.00)
-4.00 (0.00)
-3.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.18 (0.38)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.49 (0.16)
-2.00 (0.00)
-1.67 (0.00)
-1.67 (0.00)
-2.00 (0.00)
-2.08 (0.27)
-2.00 (0.00)
-1.64 (0.28)
-1.65 (0.28)
-25
-CART-depth-2
-CART-size-4
-ConTree-2
-HODT
-Ground Truth
-89.76 (4.47)
-93.58 (4.13)
-92.29 (3.65)
-98.35 (1.00)
-97.67 (0.91)
-87.13 (4.93)
-90.28 (4.50)
-89.55 (4.17)
-97.62 (1.20)
-100 (0.00)
-2.98 (0.14)
-4.00 (0.00)
-2.98 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-3.24 (0.43)
-2.00 (0.00)
-2.00 (0.00)
-2.00 (0.00)
-1.99 (0.046)
-2.53 (0.17)
-1.99 (0.05)
-1.67 (0.00)
-1.67 (0.00)
-1.99 (0.09)
-2.05 (0.35)
-1.99 (0.05)
-1.67 (0.27)
-1.65 (0.28)
-Figure 11: The effect of noise on training data. Training data size = 100. Ground truth trees are size 2.
+0 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.29 (4.74) 94.20 (3.70) 92.73 (3.90) 99.88 (0.26) 100 (0.00) 87.23 (5.14) 90.32 (4.24) 89.55 (4.10) 98.08 (1.00) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.2 (0.399) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.065) 2.51 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.09 (0.33) 2.00 (0.00) 1.63 (0.29) 1.63 (0.28)
+5 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.16 (4.63) 93.98 (3.78) 92.62 (3.79) 99.62 (0.40) 99.59 (0.44) 87.16 (5.06) 90.32 (4.20) 89.50 (3.94) 98.10 (1.01) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.05) 2.53 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.01) 2.05 (0.32) 2.00 (0.00) 1.64 (0.28) 1.65 (0.28)
+10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 90.30 (4.55) 93.98 (3.57) 92.50 (3.65) 99.07 (0.62) 98.81 (0.71) 87.36 (5.09) 90.30 (4.55) 89.38 (4.09) 97.50 (0.90) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.28 (0.45) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.065) 2.54 (0.17) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.03 (0.34) 2.00 (0.00) 1.58 (0.29) 1.65 (0.28)
+10 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.86 (4.91) 93.78 (3.92) 92.52 (3.83) 99.05 (0.66) 98.72 (0.80) 87.06 (4.97) 90.27 (4.42) 89.31 (4.38) 97.90 (1.10) 100 (0.00) 2.96 (0.20) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.32 (0.47) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.07) 2.55 (0.18) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.02) 2.03 (0.33) 2.00 (0.00) 1.63 (0.28) 1.65 (0.28)
+20 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.75 (4.98) 93.75 (3.70) 92.39 (3.72) 98.65 (0.87) 98.11 (1.04) 86.61 (5.57) 90.21 (4.27) 89.25 (4.05) 97.40 (11.50) 100 (0.00) 3.00 (0.00) 4.00 (0.00) 3.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.18 (0.38) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 2.49 (0.16) 2.00 (0.00) 1.67 (0.00) 1.67 (0.00) 2.00 (0.00) 2.08 (0.27) 2.00 (0.00) 1.64 (0.28) 1.65 (0.28)
+25 CART-depth-2 CART-size-4 ConTree-2 HODT Ground Truth 89.76 (4.47) 93.58 (4.13) 92.29 (3.65) 98.35 (1.00) 97.67 (0.91) 87.13 (4.93) 90.28 (4.50) 89.55 (4.17) 97.62 (1.20) 100 (0.00) 2.98 (0.14) 4.00 (0.00) 2.98 (0.00) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 3.24 (0.43) 2.00 (0.00) 2.00 (0.00) 2.00 (0.00) 1.99 (0.046) 2.53 (0.17) 1.99 (0.05) 1.67 (0.00) 1.67 (0.00) 1.99 (0.09) 2.05 (0.35) 1.99 (0.05) 1.67 (0.27) 1.65 (0.28)
+Table 5: The effect of noise on training data. Training data size = 100. Ground truth trees are size 2.
 Finally, in the last set of experiments, we examined the effect of noise in the features of the data.
@@
 haberman 283 3 2 75.58/73.68 (1.62/3.14) 76.991/72.98 (0.93/3.06) 74.25/72.63 (1.26/4.52) 75.31/73.68 (1.82/2.94) 79.38/78.25 (1.23/3.60) 80.70/78.24 (1.39/4.87)
-BldTrns 502 4 2 75.71/72.08 (1.05/2.02 77.16/71.09 (0.60/01.70) 75.711/72.08 (1.05/02.02) 76.91/74.06 (0.56/03.09) 79.38/78.50 (1.16/4.37) 80.71/78.25 (1.12/3.86)
+BldTrns 502 4 2 75.71/72.08 (1.05/2.02) 77.16/71.09 (0.60/01.70) 75.711/72.08 (1.05/02.02) 76.91/74.06 (0.56/03.09) 79.38/78.50 (1.16/4.37) 80.71/78.25 (1.12/3.86)
 spesis 975 3 2 94.13/93.64 (0.30/1.06) 94.26/93.74 (0.30/0.88) 94.13/93.64 (0.30/1.06) 94.18/93.64 (0.35/1.06) 94.97/93.85 (0.44/1.95) 95.39/93.85 (0.43/1.95)
@@
 GlsId 213 9 6 62.35/62.33 (1.18/5.38) 67.29/62.79 (1.15/04.88) 61.76/62.33 (1.18/5.38) 66.47/63.72 (1.12/5.22) 71.18/61.86 (2.46/4.53) 74.25/62.79 (2.29/5.20)
-Parkinsons 195 22 2 87.95/84.62 (1.48/5.85) 93.20/82.56 (1.04/5.23) 87.82/85.13 (1.67/6.36 90.77/86.67 (1.04/5.71 93.72/89.74 (1.66/3.14) 94.49/89.74 (2.15/3.14
+Parkinsons 195 22 2 87.95/84.62 (1.48/5.85) 93.20/82.56 (1.04/5.23) 87.82/85.13 (1.67/6.36) 90.77/86.67 (1.04/5.71) 93.72/89.74 (1.66/3.14) 94.49/89.74 (2.15/3.14
 Diabetic 1146 19 2 65.55/64.00 (0.92/3.74) 67.36/63.48 (0.79/3.95) 65.55/64.00 (0.92/3.74) 67.53/62.96 (1.05/2.66) 79.93/76.52 (1.17/1.71) 80.09/76.35 (1.06/2.38)
@@
 AucVer 2043 7 2 90.42/89.58 (0.55/1.57) 91.00/89.98 (0.40/1.59) 90.27/89.49 (0.40/1.61) 90.42/89.58 (0.55/1.57) 89.65/88.46 (0.35/1.99) 89.87/88.66 (0.34/1.83)
-Ai4iMF 10000 6 2 97.25/97.07 (0.11/0.36) 97.39/97.31 0.07/0.26) 97.08/97.00 00.27/00.34) 97.16/97.03 0.22/0.36) 97.79/97.95 0.11/0.18) 97.90/98.00 0.12/0.15)
+Ai4iMF 10000 6 2 97.25/97.07 (0.11/0.36) 97.39/97.31 (0.07/0.26) 97.08/97.00 (00.27/00.34) 97.16/97.03 (0.22/0.36) 97.79/97.95 (0.11/0.18) 97.90/98.00 (0.12/0.15)
 VoicePath 704 2 2 96.87/95.60 (0.51/2.12) 97.16/96.31 (0.56/2.43) 96.87/95.60 (0.51/2.12) 97.16/95.04 (0.48/1.85) 97.69/97.44 (0.28/1.08) 97.90/97.31 (0.32/1.17)
@@
 Overall, our findings indicate that HODT not only achieves higher accuracy but also offers a more general and robust framework for decision tree construction, paving the way for broader applications in machine learning where interpretability and predictive power are equally critical.
-6 Conclusion and future directionsa
+6 Conclusion and future directions
 Conclusion
@@
 There remain several limitations in our current research that warrant further investigation.
-First, implementation of additional algorithms.
-In our experiments, we implemented only $\mathit{odt}_{\text{rec}}$ and $\mathit{odt}_{\text{vec}}$, while $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ were not considered.
-We note that $\mathit{odt}_{\text{depth}}$ is challenging to parallelize compared with $\mathit{odt}_{\text{size}}$ , $\mathit{odt}_{\text{kperms}}$ and shares features with $\mathit{odt}_{\text{vec}}$.
-Nevertheless, both $\mathit{odt}_{\text{depth}}$ and $\mathit{odt}_{\text{kperms}}$ has their advantages: $\mathit{odt}_{\text{depth}}$ has advantages for constructing large trees efficiently using a depth-first search strategy.
-It would be interesting to explore whether $\mathit{odt}_{\text{depth}}$ could be as effective for hyperplane decision trees as it is for axis-parallel decision trees, as reported in Brița et al.
-(2025).
-Second, improved heuristics.
-The current two heuristics become ineffective when $K$ and $D$ are large, providing only marginal improvements or even deteriorating accuracy.
-Randomly generating hyperplanes is ineffective because the probability of finding feasible nested combinations is low.
-Future work could focus on developing more powerful heuristics for constructing hyperplane decision trees in high-dimensional or large-$K$ settings.
-One possible approach is to combine depth-first search with coreset selection methods.
-Third, conflicts between the general position assumption and categorical data.
+First, conflicts between the general position assumption and categorical data.
 The correctness of $\mathit{hodt}$ relies on the general position assumption for data points.
@@
 Investigating ways to incorporate this assumption into hypersurface decision tree algorithms could lead to more efficient methods for categorical datasets.
-Fourth, algorithms with mixed splitting rules.
+Second, algorithms with mixed splitting rules.
 In Part I of this paper, we proposed a generator for mixed-splitting rules.
@@
 Future work could explore the construction of potentially more flexible models using these rules.
-Finally, extensions to random forests.
-Two key results from Breiman (2001) suggest promising directions for constructing random forests with hypersurface decision trees.shows that adding more trees does not lead to overfitting but converges to a limiting generalization error, while Theorem 2.3 establishes that generalization performance improves with the strength of individual classifiers.
+Third, extensions to random forests.
+Two key results from Breiman (2001) suggest promising directions for constructing random forests with hypersurface decision trees.
+Theorem 1.2 shows that adding more trees does not lead to overfitting but converges to a limiting generalization error, while Theorem 2.3 establishes that generalization performance improves with the strength of individual classifiers.
 These insights align naturally with our framework: even in our Python implementation, the $\mathit{sodt}$ program can efficiently solve $1\times 10^{6}$ feasible nested combinations in tens of seconds.
@@
 A.1 Nested-combination generator
-  1. 1.
-    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
-  2. 2.
-    Output: Array of $\left(K,G\right)$-nested-combinations
-  3. 3.
-    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
-  4. 4.
-    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
-  5. 5.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
-  6. 6.
-    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
-  7. 7.
-    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
-  8. 8.
-      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
-  9. 9.
-      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
-  10. 10.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
-  11. 11.
-    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
-  12. 12.
-    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
-  13. 13.
-       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
-  14. 14.
-        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
-  15. 15.
-    return $\mathit{ncss}\left[K\right]$
-Algorithm 5 $\mathit{nestedCombs}$
+Algorithm 5 Nested Combination Generator ($\mathit{nestedCombs}(K,G,\mathit{xs})$)
+1:Input: $\mathit{xs}$: data list, length NN; $K$: outer combination size; $G$: inner-combination size
+2:Output: Array of $(K,G)$-nested-combinations
+[…]
+[…]
+[…]
+6:for $n\in\mathit{range}(0,N)$ do
+7:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
+8:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
+9:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
+10:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
+11:  end for
+12:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n+1}{G}$
+13:  for $i\in\mathit{range}(C_{1},C_{2})$ do
+14:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
+15:    $\mathit{ncss}[k]\leftarrow\mathit{map}(\cup[i],\mathit{ncss}[k-1])\cup\mathit{ncss}[k]$ $\triangleright$ Update nested combinations
+16:   end for
+17:  end for
+18:end for
+19:return $\mathit{ncss}[K]$
 A.2 Incremental ancestry relation matrix generator
-  1. 1.
-    Input: $\mathit{xs}$: input data list of length $N$; $K$: Outer combination size (nested combination); $G$: inner-combination size (ordinary combination), determined by polynomial degree $M$ of hypersurface
-  2. 2.
-    Output: Array of $\left(K,G\right)$-nested-combinations
-  3. 3.
-    $\mathit{css}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize combinations
-  4. 4.
-    $\mathit{ncss}=\left[\left[\left[\>\right]\right],\left[\right]^{k}\right]$ // initialize nested-combinations
-  5. 5.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{empty}\left(\left(\begin{array}[]{c}N\\ D\end{array}\right),N\right)$ // initialize $\mathit{asgn}^{+},\mathit{asgn}^{-}$ as two empty $\left(\begin{array}[]{c}N\\ D\end{array}\right)\times N$ matrix
-  6. 6.
-    for $n\leftarrow\mathit{range}\left(0,N\right)$ do: $\mathit{range}\left(0,N\right)=\left[0,1,\ldots,N-1\right]$
-  7. 7.
-    for $j\leftarrow\mathit{reverse}\left(\mathit{range}\left(G,n+1\right)\right)$ do:
-  8. 8.
-      $\mathit{updates}=\mathit{reverse}\left(\mathit{map}\left(\cup\rho_{M}\left(\mathit{xs}\right)\left[n\right],\mathit{css}\left[j-1\right]\right)\right)$
-  9. 9.
-      $\mathit{css}\left[j\right]=\mathit{css}\left[j\right]\cup\mathit{updaets}$ // update $\mathit{css}$ to generate combinations in revolving door ordering,
-  10. 10.
-    $\mathit{asgn}^{+},\mathit{asgn}^{-}=\mathit{genModels}\left(\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // use $G$-combination to generate the positive prediction and negative prediction of each hyperplanes and stored in $\mathit{asgn}^{+},\mathit{asgn}^{-}$
-  11. 11.
-    $C_{1}=\left(\begin{array}[]{c}n\\ G\end{array}\right)$, $C_{2}=\left(\begin{array}[]{c}n+1\\ G\end{array}\right)$
-  12. 12.
-    for $i\leftarrow\mathit{range}\left(C_{1},C_{2}\right)$ do:
-  13. 13.
-       for $k\leftarrow\mathit{reverse}\left(\mathit{range}\left(K,i+1\right)\right)$ do:
-  14. 14.
-        $\mathit{ncss}\left[k\right]=\mathit{map}\left(\cup\left[i\right],\mathit{ncss}\left[k-1\right]\right)\cup\mathit{ncss}\left[k\right]$ // update nested combinations
-  15. 15.
-        $\mathit{ncss}\left[k\right]=\mathit{updates}_{\text{arMat}}\left(i,\mathit{ncss}\left[k\right],\mathit{css}\left[G\right],\mathit{asgn}^{+},\mathit{asgn}^{-}\right)$ // $\mathit{update}_{\text{arMat}}$ update $\left(k,G\right)$-nested-combinations and its associated ancestry relation matrix, defined in Algorithm 1 for hyperplane spliting rules.
-  16. 16.
-    return $\mathit{ncss}\left[K\right]$
-Algorithm 6 $\mathit{\mathit{\mathit{nestedCombsFA}}}\left(K,G,\mathit{xs}\right)$
-A.3 Experiments for $K>3$ decision tree
+Algorithm 6 Nested Combination Generator with Ancestry ($\mathit{nestedCombsFA}(K,G,\mathit{xs})$)
+1:Input: $\mathit{xs}$: data list, length NN; $K$: outer combination size; $G$: inner-combination size
+2:Output: Array of $(K,G)$-nested-combinations
+[…]
+[…]
+[…]
+6:for $n\in\mathit{range}(0,N)$ do
+7:  for $j\in\mathit{reverse}(\mathit{range}(G,n+1))$ do
+8:   $\mathit{updates}\leftarrow\mathit{reverse}(\mathit{map}(\cup\rho_{M}(\mathit{xs})[n],\mathit{css}[j-1]))$
+9:   $\mathit{css}[j]\leftarrow\mathit{css}[j]\cup\mathit{updates}$ $\triangleright$ Update combinations
+10:   $\mathit{asgn}^{+},\mathit{asgn}^{-}\leftarrow\mathit{genModels}(\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Generate predictions
+11:  end for
+12:  $C_{1}\leftarrow\binom{n}{G}$, $C_{2}\leftarrow\binom{n+1}{G}$
+13:  for $i\in\mathit{range}(C_{1},C_{2})$ do
+14:   for $k\in\mathit{reverse}(\mathit{range}(K,i+1))$ do
+15:    $\mathit{ncss}[k]\leftarrow\mathit{map}(\cup[i],\mathit{ncss}[k-1])\cup\mathit{ncss}[k]$ $\triangleright$ Update nested combinations
+16:    $\mathit{ncss}[k]\leftarrow\mathit{updates}_{arMat}(i,\mathit{ncss}[k],\mathit{css}[G],\mathit{asgn}^{+},\mathit{asgn}^{-})$ $\triangleright$ Update ancestry matrix, Alg. 1
+17:   end for
+18:  end for
+19:end for
+20:return $\mathit{ncss}[K]$
+A.3 Experiments for $K>3$ decision trees
+The experimental results for $K>3$ decision trees are given in Table 7.
 Dataset NN $D$ $C$ CART-depth $d=3$ ConTree $d=3$ CART-size $K=4$ CART-size $K=5$ CART-size $K=6$ HODT $K=4$ HODT $K=5$ HODT $K=6$